### v1

# Evaluating Diversity in Automatic Poetry Generation

Abstract
Natural Language Generation (NLG), and more generally generative AI, are among the currently most impactful research fields. Creative NLG, such as automatic poetry generation, is a fascinating niche in this area. While most previous research has focused on forms of the Turing test when evaluating automatic poetry generation — can humans distinguish between automatic and human generated poetry — we evaluate the diversity of automatically generated poetry, by comparing distributions of generated poetry to distributions of human poetry along structural, lexical, semantic and stylistic dimensions, assessing different model types (word vs. character-level, general purpose LLMs vs. poetry-specific models) and types of fine-tuning (conditioned vs. unconditioned). We find that current automatic poetry systems are considerably underdiverse along all dimensions — they tend to memorize, do not rhyme sufficiently, are semantically too uniform and even do not match the length distribution of human poetry. Among all models explored, character-level style-conditioned models perform slightly better. Our identified limitations may serve as the basis for more genuinely creative future poetry generation models.
Evaluating Diversity in Automatic Poetry Generation
Anonymous ACL submission
1 Introduction
A key aspect of creative language generation is the ability to create new, original and interesting text, cf. colton2012computational; gatt2018survey; yi2020mixpoet; Elgammal2017CANCA. To date, extremely little attention has been given to the evaluation of originality and creativity in recent creative text generation models such as those for automatic poetry generation, despite renewed interest in the context of recent LLMs franceschelli2023creativity. In fact, existing automatic poetry generation models are typically not evaluated regarding how different generated poems are from existing poems in the training set but with the Turing test: can humans distinguish whether a poem is human authored or automatically generated (hopkins-kiela-2017-automatically; lau2018deep; Manjavacas2019ARS)? However, this form of Turing test and other similar forms of human evaluation may contain an overlooked risk of failure: namely, if the automatically generated instances are (near-)copies of training data instances.
In this work, we fill this gap and evaluate, for the first time, automatic poetry generation systems for their diversity. As human evaluation is generally not well suited to assess diversity hashimoto-etal-2019-unifying, we automatically measure diversity by comparing distributions of generated and existing poems along formal, semantic and stylistic dimensions. This yields much better evidence of the models’ creative capabilities in contrast to being mere ‘stochastic parrots’.
Our main contributions are: (i) we conceptualize diversity of poetry generation systems along different dimensions: diversity on the structural, lexical, semantic and stylistic level; (ii) we assess different types of automatic poetry generation systems for diversity: general purpose word and character-level LLMs, both unconditioned and style-conditioned ones, on the one hand, and poetry-specific models, on the other hand; (iii) we evaluate each class of model for diversity across the different dimensions, by comparing the distribution of the human authored training data set to the distribution of generated poems. We find that on a distributional level, generated poems are considerably different from human ones. Concerning general purpose LLMs, some of them exhibit very high risk of memorization — an extreme form of lack of diversity — and this depends on the size of the training data set, the size and type of the LLM, and the type of training, as we show. Character-level style-conditioned general-purpose LLMs are most diverse.
Our work prepares the groundwork for truly creative generative AI models (veale2020leaps) and also has implications for the detection of generative AI (Sadasivan2023CanAT).
2 Related Work
Our work connects to research on diversity and automatic poetry generation, which we now discuss.
Diversity
Building systems able to generate diverse output has been a long-standing concern in NLG research (reiter-sripada-2002-squibs; van-deemter-etal-2005-squibs; foster-white-2007-avoiding) and remains a central issue in neural NLG (holtzman2019curious). The need for careful analysis of NLG systems diversity – beyond an assessment of the quality or fluency of single-best generation outputs – has been widely acknowledged (gatt2018survey; hashimoto-etal-2019-unifying; mahamood-zembrzuski-2019-hotel; celikyilmaz2020evaluation; tevet-berant-2021-evaluating; schuz-etal-2021-diversity). A well-known finding from this line of research is that neural NLG systems typically face a quality-diversity trade-off (ippolito-etal-2019-comparison; caccia2020; wiher2022decoding): their outputs are either well-formed and fluent or diverse and variable.
Work on evaluating diversity of NLG typically uses automatic metrics that quantify to what extent different outputs by the same system vary (hashimoto-etal-2019-unifying). In practice, though, evaluations of diversity in NLG differ widely across tasks (tevet-berant-2021-evaluating) and even adopt different notions of diversity (zarriess-etal-2021-decoding). At the same time, most of these notions focus on lexical or semantic aspects of diversity, e.g., local lexical diversity. For instance, ippolito-etal-2019-comparison compare decoding methods in dialog generation and image captioning, assessing lexical overlaps in -best NLG outputs for the same input. Global lexical diversity, on the other hand, measures whether the NLG system generates different outputs for different inputs. For instance, van-miltenburg-etal-2018-measuring define the global diversity of image captioning systems as their ability to generate different captions for a set of inputs, using metrics like the number of types in the output vocabulary, type-token ratio, and the percentage of novel descriptions. Similarly, hashimoto-etal-2019-unifying view diversity as related to the model’s ability to generalize beyond the training set, i.e., generate novel sentences.
Besides lexical diversity, work on open-ended or creative text generation tasks has been interested in diversity at a more general semantic level. For instance, zhang2018 and stasaski-hearst-2022-semantic aim at building dialogue systems that generate entertaining and semantically diverse responses in chit-chat dialog, where the goal is to avoid “safe and bland” responses that “average out” the sentences observed in the training set. Here, semantic diversity has been measured, e.g., with the help of embedding-based similarity (du-black-2019-boosting).
In our work on diversity in poetry generation, we complement these lexical and semantic aspects of diversity with aspects of formal diversity. We thus explore whether automatic poetry generation systems are able to capture the ‘full bandwidth’ of realizations of poetry found in the data distribution with which they have been trained, focusing mostly on global diversity.
Poetry generation
Automatic poetry generation is a long standing dream of AI research, dating back at least to the mid 20th century (e.g., Theo Lutz’ Stochastische Texte). While early modern systems were heavily hand-engineered gervas2001expert, more recent approaches are all trained on collections of human poetry lau2018deep; jhamtani2019learning; agarwal-kann-2020-acrostic but still extensively utilize human guidance e.g. to enforce formal characteristics of poetry such as rhyming wockener-etal-2021-end. belouadi2022bygpt5 have recently released a character-level decoder-only LLM (ByGPT5) capable of learning style-constraints such as rhyming without human involvement in model design.
In our work, we explore varying poetry generation models with regard to diversity: poetry-specific models that use hand-engineered architectures as well as general purpose LLMs, including ByGPT5.
3 Diversity In Poetry Generation
We first conceptualize diversity in poetry generation using formal and semantic criteria. As our dataset, we use QuaTrain (belouadi2022bygpt5) consisting of quatrains (in English and German). We describe it in more detail in §5.1 below.
Memorization.
In poetry, as in other forms of art, creativity (sternberg1999handbook) plays a central role. A basic aspect of creativity is the models’ ability to generate poems that are different from the training data, i.e. have not been memorized as a whole. Therefore, we consider a low or minimal degree of memorization as a pre-requisite for diversity and analyze the portion of generated poems that are (near-)copies from the training data. To examine memorization, we proceed as in belouadi2022bygpt5. We apply the Ratcliff-Obershelp similarity (ratcliff1988pattern) to compare each poem in a sample with poems in the training corpus. If a generated quatrain exhibits a similarity score of 0.7 with a quatrain in the training data, we classify it as memorized. A quatrain can be divided into 4 verses or 2 couplets; thus, we also inspect memorization at the verse and couplet levels by comparing each verse or couplet in a sample to those in the training data. Higher thresholds for classification are used for these finer-grained comparison levels, as shorter texts have higher chances of being more similar in general. 11todo: 1YC: maybe add a table in the appendix to demonstrate the couplets/verses classified as memorized with threshold 0.7. Specifically, a verse with a similarity score 0.9 or a couplet 0.8 is considered as memorized. We define the memorization score of a sample as the proportion of memorized quatrains in that sample. How much LLMs memorize from their training data has been a question of central concern recently mccoy-etal-2023-much.
Poem length.
Within a sample of generated poems, we consider differences at the level of poem length, i.e., their number of tokens, as a basic aspect of diversity at the formal or structural level. We analyze to what extent the length distribution of generated poems differs from the distribution in the training data. We define the length of a quatrain as the number of tokens contained: we eliminate all punctuation symbols and split the remaining text by white space. We report mean length, standard deviation, minimal and maximal length of samples. We additionally deploy distance measures between training data distribution and generated samples, in particular, a metric called histogram intersection (swain1991color), which measures the intersection area of two normalized histograms (and therefore returns values between 0 and 1).
Rhyme patterns.
As a more complex dimension of formal diversity, we consider rhyming as a central aspect that characterizes the structure of a poem. Diversity can then be assessed by comparing rhyme distributions between generated samples and training data. In order to classify rhymes in our samples, we use the same classifier used to annotate QuaTrain. We distinguish between true rhymes, which involve different words, and repetitions, which refer to rhymes based on the same word.
Lexical diversity.
Lexical diversity is a standard aspect of diversity evaluation in NLG and is used to assess how generation outputs vary in their vocabulary, either at the local text level or at the global corpus level. We use the following metrics to measure the lexical diversity for both the training data and the generated samples: (i) Averaged type token ratio (ATTR). We calculate ATTR as the average of all type token ratios (richards1987type) (TTRs) for each quatrain in a sample, i.e. as a measure of local lexical diversity. (ii) Moving average type token ratio (MATTR). The MATTR (covington2010cutting) acts on the corpus level and calculates a moving average by sliding through the corpus using a window of fixed size. We deploy this metric as a measure of global lexical diversity. (iii) Measure of textual, lexical diversity (MTLD). The MTLD (mccarthy2005assessment) is calculated as the average length of a substring that maintains a specified TTR level. MTLD is deployed to measure lexical diversity on a global scale.
Semantic diversity.
Even if a poetry generation system does not directly copy data from the training data, the generated poems may still be semantically very similar to the training data distribution. We employ a multilingual distilled version of Sentence-BERT (SBERT) reimers-gurevych-2019-sentence as dense vector representations to measure semantic similarity between poems: (i) across the human train set and the generated poems, (ii) within human and generated poems. In particular, for each generated quatrain, we note down the similarity value of the most similar human quatrain, then report the average over all those maximum similarity values. We proceed analogously within the human training data and within the automatically generated poems.
4 Models
Our experiments use 2 different model classes.
4.1 Poetry-specific models
Deepspeare.
Deepspeare lau2018deep is specifically designed for poetry generation. Its core architecture consists of an LSTM language model, a pentameter model (specifically designed to learn iambic meter) and a rhyme model. During training, it takes sonnets as input data (three quatrains followed by a couplet) but ultimately processes the contained quatrains by splitting any given sonnet. The rhyme model processes ending words of quatrain verses and uses a margin-based loss to discriminate between rhyming and non-rhyming words. It is not limited to specific rhyme patterns but assumes that rhymes exist in the data. At inference time, Deepspeare generates quatrains.
Structured Adversary.
Like Deepspeare, Structured Adversary (SA) jhamtani2019learning incorporates different components: an LSTM language model and a discriminator used to decide whether line endings are typical for poetry. Both components are organized in an adversarial setup, where the language model acts as a generator, trying to generate poems that are misclassified by the discriminator, while the discriminator is trained to distinguish generated poems from real ones. SA is trained with sonnets as input data. At inference time, it generates quatrains.
4.2 General purpose LLMs
All models in this category are decoder-only transformer architectures. In our experiments, we train them in an unconditioned and style-conditioned manner (see Section 5.2).
GPT2
GPT2 radford2019language is the last GPT model made publicly available. It is a large word level transformer-based language model pretrained on approximately 40 GB of text. Four different model versions were released, with the number of parameters ranging from 125 million to 1.5 billion for the largest. In this work, we deploy two model versions: GPT2-small (125M parameters) and GPT2-large (774M parameters) for both English and German.
GPTneo
GPTneo (black-etal-2022-gpt) is an open-source token level LLM by EleutherAI (https://www.eleuther.ai/) with the aim to provide publicly available replications of GPT3. It is pre-trained on 825 GB of text data. Currently, four versions have been released, with the number of parameters ranging from 125 million up to 20 billion. We deploy GPTneo-small and GPTneo-xl with 125M and 1.3B parameters for English. GPTneo is not available for German.
LLaMA2/3
ByGPT5
ByGPT5 (belouadi2022bygpt5) is a decoder-only character level LLM based on the encoder-decoder character level model byT5 (xue-etal-2022-byt5) where the encoder part of byT5 is completely removed, reducing the number of parameters by 75%. The remaining decoder-only model is then pretrained using OpenWebText for English (38GB text data) and CC100 (conneau-etal-2020-unsupervised) (67GB text data) for German. Three versions are released for both English and German, with model sizes ranging from 73 to 298M parameters. We use ByGPT5-base (140M params) and ByGPT5-medium (290M) for both English and German.
5 Experimental Setup
5.1 Training Data
DE | EN | |||
QuaTrain | SonNet | QuaTrain | SonNet | |
Train | 253,843 | 72,526 | 181,670 | 51,905 |
Dev | 28,205 | 8,058 | 20,186 | 5,767 |
Total | 282,048 | 80,584 | 201,856 | 57,672 |
We use QuaTrain, a large dataset of quatrains published by belouadi2022bygpt5. It consists of English and German quatrains and has been generated by aggregating different publicly available poetry datasets. QuaTrain contains human written quatrains but mixes them synthetically: every sequence of four consecutive lines from the underlying human data are included in order to increase dataset size. QuaTrain is automatically annotated for meter and rhyme using high-quality classifers (especially for rhyme). 22todo: 2YC: @Hannes, could you please add here how you preprocessed the data to drop the redundant ones? H: on it! Because our focus lies on the diversity of model outputs, we have to avoid repetitions in the training data created by the data augmentation methods used in its creation. To avoid lines appearing multiple times, we first parse the dataset sequentially, eliminating quatrains that overlap the preceding one. Because this method does not eliminate all overlaps, we then use an heuristic, deleting the ten percent of the quatrains which have the biggest overlap with other quatrains until there is no overlap remaining.
We use a ratio of 9:1 to split the data into train and dev sets; we do not keep a test set since no held-out human data is needed for generation or evaluation. Further, as DeepSpeare and SA are designed to process sonnets and/or limerick data, we create pseudo sonnets for them using the corresponding data splits. Specifically, for each sonnet, we randomly draw three quatrains and one couplet from the QuaTrain datasets, ensuring that each comes from a different original quatrain. Table 1 provides the resulting data statistics.
5.2 Training
DeepSpeare.
DeepSpeare (lau2018deep) leverages pretrained static word vectors. We use QuaTrain and SonNet to train our own Word2vec embeddings (mikolov2013efficient) and the final sonnet models respectively. For the sonnet model training, we use a batch size of 128 and apply early stopping with a patience of 5 epochs; default settings are maintained for the other hyperparameters.
SA.
We use the same word vectors and training data splits as for DeepSpeare. Training SA involves 1) pretraining the discriminator’s encoder using a publicly available pronouncing dictionary ; 2) training the LM component; 3) training a final aggregated model in a generative adversarial setup. We train the discriminators with a batch size of 128, the LMs with a batch size of 64, and the final sonnet models with a batch size of 128; here, we also implement early stopping with a patience of 5 epochs.
Unconditioned LLMs.
In this setup, we fine-tune our decoder-only LLMs in an unconditioned manner: we process quatrains during training without passing any information about rhyme (or meter). All models except GPTneo (being available only in English) are trained both in English and German. We fine-tune all models (English and German) for 50 epochs, with early stopping using a patience of 3 epochs.
Style-conditioned LLMs.
In contrast to unconditioned training, we provide information about rhyme (and meter) by prepending special style tokens to each quatrain during training. This follows the setup of belouadi2022bygpt5 and makes models explicitly aware of different rhyme schemes. As for the unconditioned variants, all models except GPTneo are trained in English and German. We use the same hyperparameter settings as for the unconditioned trained LLMs.
Summary.
We end up with 36 models that can be assigned to three categories: 1) Poetry specific LSTM-based models (DeepSpeare and SA). Besides a language model part, these models incorporate additional specialized components to handle poetry-specific stylistic features such as rhyme. We have 2 models for each language. 2) Unconditioned LLMs (transformer-based decoder-only general purpose LLMs). These models do not possess any specialized architecture for poetry. No information about meter or rhyme has actively been passed during training. We have two subcategories: word and character level models. The first group (the GPT and LLaMA series) processes data on the word/subword level. ByGPT5 represents the character-level group. We have 9 models for English and 7 for German. 3) Style-conditioned LLMs. These have the same architecture, models, and subgroups as category 2. Information about rhyme (and meter) is passed in the form of special tokens during training (only). We use the following notation to distinguish the size and training manners of the LLMs: a subscript of S/L indicates whether it is a smaller/larger version, and a superscript of “con” stands for conditioned training. E.g., GPT2 and GPT2 represent the unconditioned and conditioned trained GPT2 small models, respectively.
Table 4 (appendix) provides an overview of all models belonging to the second and third category (transformer-based LLMs).
5.3 Sampling
From each model class, we randomly draw 1000 generated poems. Whenever wo do a direct comparison between training and generated data (e.g. when comparing lexical diversity), we randomly draw 10 samples of size 1000 (matching the sample size) from the train set and use mean results as representatives. We deploy this strategy to mitigate the large discrepancy in size between human data and generated poems. For all LLMs, we employ consistent decoding strategies for generation: we use the default settings of the LLaMA2 chat models on Hugging Face111https://huggingface.co/spaces/huggingface-projects/llama-2-7b-chat but limit the number of newly generated tokens to 100 for the token-level models or 300 for the character-level ByGPT5 models.44todo: 4YC: @Hannes, do you have different sampling strategies for bygpt5? H: Other than the maximum length it’s the same. In Section 7, we briefly discuss the effects of various decoding strategies.
6 Experiments and Results
We first investigate structural properties of the generated poems (repetition of instances on a surface level, length distributions, rhyming), then consider lexical and semantic properties.
Memorization
On quatrain-level, no model exhibits memorization, with all obtaining a memorization rate of 0.
Considering couplet-level memorization, 23 out of 36 models show zero memorization, while 13 models display scores between 0.05% and 0.15%. The poetry-specific models, SA and DeepSpeare, as well as the character-level ByGPT5 models, exhibit no memorization; in contrast, GPT2 and GPTNeo models show the highest rates on average (up to 0.15% for German and 0.10% for English). When comparing models of the same architecture and training methods but varying sizes, differences are found in 6 out of 14 cases. In 5 cases, larger models have 0.05%-0.10% higher absolute memorization scores than their smaller counterparts (the English GPT2con and LLaMA2con models, and the German GPT2con, GPTNeocon, LLaMA2 models); the only exception is the English GPTNeo models, where the smaller one has a 0.05% higher memorization rate. On the other hand, there is no distinct boundary between unconditioned and conditioned trained models: in the 6 cases where discrepancies in memorization rates exist, conditioned models outperform their unconditioned ones three times, and vice versa, with absolute differences of 0.05%-0.1%.
In the verse-level evaluation, the poetry-specific models perform best overall (0.4%-0.83% for German and 0.1%-0.83% for English), followed by the ByGPT5 models (0.68%-1.3% for German and 0.58%-1.23% for English). Again, GPT2 is worst for German, exhibiting memorization rates of 4.38%-8.7%, whereas, for English, GPTNeo exhibits the highest rates, ranging from 3.5%-5.6%. Concerning different model sizes, we again see that larger models memorize more than their smaller counterparts: in 9 out of 14 cases, larger models show higher memorization rates, with an average absolute increase of 0.15%. Furthermore, unlike in the couplet-level evaluation, here, each unconditioned model exhibits a strictly higher memorization rate compared to its conditioned counterpart, with an absolute increase of 1.47% on average.
In summary, with respect to memorization: (1) Among model families, the poetry-specific and character-level models are the most diverse. (2) Larger models are less diverse compared to their smaller versions. (3) Conditional training enhances model diversity.
Length
Table 5 (appendix) reports statistics on the length of poems, both human and automatically generated. The mean length of human written poems is 28 in English and 24 in German. Histogram intersection values between samples generated by the models and the human written data range from 0.61 to 0.88 in German (LLaMA2 and SA) and from 0.48 to 0.92 in English (GPTNeo and SA). While the SA models fit the distribution of the human written poems the best, the character-level ByGPT5 models also perform well consistently with histogram intersection values between 0.77 and 0.85. The poems generated by German LLaMA2 and English GPTNeo are too short and not diverse enough (in terms of standard deviation). The poetry-specific DeepSpeare models do not match the human distribution very well either, with intersection values of 0.63 and 0.57 for German and English respectively. Here too, poems lengths are too small and not diverse enough. Conditioned models seem to fit the training data better across the board, the only exceptions being German ByGPT5 and English LLaMA2. Figure 1 illustrates the length distribution of human written poems, SA and GPTNeo for English.
Rhyme
Figure 2 (a) shows the distributions of rhyme schemes in our human training datasets for German; Figure 3 (a) shows the corresponding distribution for English. 55todo: 5SE: Most rhymes in the training data are classified as real rhymes. For both languages, less than 15% of all quatrains in training do not rhyme at all (rhyme scheme ABCD). Excluding ABCD, the top 3 dominant rhyme schemes by appearance are ABAB, AABB and ABCB for both datasets, with a total share of approximately 60% in each language. German has a higher proportion of ABAB (above 35%), while English has ABAB and AABB in roughly equal proportions (25%). In Table 6 (appendix), we report the entropy of all rhyme distributions as well the distance between the human distribution and model distributions, measured in cross-entropy and KL divergence. The best, worst and an average model, in terms of cross-entropy, are shown in Figures 2 and 3.
Poetry-specific models: Figure 5 (appendix) shows the distributional plots for Deepspeare and SA. We observe that DeepSpeare has a very low ratio of ABCD, considerably lower than human poems (less than 5% for both languages). The three dominating patterns are AABB, ABAB, and ABBA which (only) partially agrees with the dominating patterns in the human data. Nonetheless, DeepSpeare has the best fit of all models in terms of KL divergence and almost the best fit in terms of cross-entropy. SA has a much worse fit and produces considerably too many ABCD patterns (close to or above 30% in both languages). It has one of the worst fits to the human rhyme distributions across all models.
Figures 6 and 7 (appendix) show the distributions of rhyme patterns for unconditioned LLMs. Except for LLAMA3, all models of this kind have a high distribution of ABCD and consequently a high likelihood of producing non-rhyming poems. Thus, they have the worst fit to the human distribution, on average, among all model classes considered.
Style-conditioned LLMs are shown in Figures 8 and 9 (appendix). In general, this model class matches the human distribution closest in terms of KL divergence and cross-entropy. LLAMA3 has a particularly good fit in English, but it is noteworthy that it produces too many ABAB rhyme schemes, which may indicate a skewness due to its pre-training data. No model produces a lot of AABB rhyme pattern which abound in our human training data. Across all models in this class, the fit to the human data is still mediocre at best.
Overall, most models have clearly higher ABCD rhyming schemes than the human data, thus are underdiverse concerning rhyming. The best model class are style-conditioned LLMs (especially LLAMA3), however the poetry-specific DeepSpeare model can be considered the best individual model in terms of matching the human rhyme distribution. The character-level ByGPT5 models perform worse than word-level models without style-conditioning, but with style-conditioning they outperform the word-level models in terms of match with human rhyme distribution.
Model | ATTR (%) | MATTR (%) | MTLD |
HUMAN | 91.6 / 87.7 | 90.6 / 87.3 | 283.1 / 183.4 |
DeepSpeare | 92.6 / 89.1 | 87.9 / 84.8 | 110.0 / 89.7 |
SA | 93.0 / 88.9 | 91.0 / 87.8 | 215.6 / 162.2 |
ByGPT5 | 89.7 / 81.5 | 86.9 / 79.7 | 135.4 / 66.5 |
ByGPT5 | 91.2 / 82.5 | 88.1 / 80.5 | 151.6 / 69.9 |
GPT2 | 86.2 / 79.4 | 81.2 / 76.4 | 64.1 / 46.0 |
GPT2 | 94.2 / 87.6 | 89.5 / 83.5 | 131.8 / 81.6 |
GPTNeo | - / 78.3 | - / 74.9 | - / 40.1 |
GPTNeo | - / 86.8 | - / 81.3 | - / 61.7 |
LLaMA2 | 92.8 / 89.6 | 87.7 / 86.8 | 120.7 / 106.8 |
LLaMA2 | 94.8 / 90.2 | 90.2 / 85.7 | 150.1 / 96.0 |
LLaMA3 | 94.4 / 92.7 | 89.3 / 87.4 | 128.0 / 108.1 |
ByGPT5 | 92.2 / 85.1 | 89.5 / 83.1 | 187.1 / 94.6 |
ByGPT5 | 93.0 / 85.9 | 90.0 / 83.9 | 192.6 / 102.5 |
GPT2 | 89.2 / 84.0 | 84.2 / 81.9 | 82.0 / 70.3 |
GPT2 | 94.2 / 88.0 | 90.0 / 90.7 | 137.4 / 90.7 |
GPTNeo | - / 83.1 | - / 80.2 | - / 61.2 |
GPTNeo | - / 87.0 | - / 82.1 | - / 69.4 |
LLaMA2 | 91.1 / 90.0 | 86.8 / 88.2 | 104.4 / 109.3 |
LLaMA2 | 91.9 / 90.8 | 86.5 / 87.2 | 100.2 / 101.0 |
LLaMA3 | 93.5 / 91.7 | 89.1 / 88.3 | 128.5 / 116.3 |
Lexical Diversity.
Table 2 shows the lexical diversity results for English and German. For local diversity (ATTR), most of the models are close to the diversity in human-written poems, with the traditional models (DS, SA) and the Llama exceeding the ATTR values of human-written poems. For English, the least locally diverse poems are generated by GPT2-small, in the un/conditioned case respectively. For German, the least locally diverse models is GPT2neo-small, in the un/conditioned case respectively. The global diversity metrics (MATTR, MTLD) show different trends than ATTR, though. The MATTR metric suggests that most models do not generally achieve the level of diversity found in human poems: in English, only SA matches and slightly exceeds human diversity, in German, only the conditioned Llama3-large model exceeds human diversity. According to the MTLD metric, all models generate severely under-diverse output at the sample level. Here, the best model in English and German is , but even does not come close to the human level of global diversity. According to MTLD, style-conditioned LLMs consistently outperform their non-conditioned counterparts, with the English Llama2 models being the only exceptions here. ††todo: include barplots in the appendix? Moreover, we observe that model size affects all three lexical diversity metrics, whereby larger models are more diverse than their smaller counterparts. ††todo: produce a barplot that shows the effect of size? The effect of size is most pronounced for GPT2 in both languages, where ATTR, MATTR and MTLD substantially improve from the small to the larger model variant. It also noteworthy, though, that the more classical models, and , generally perform on par with recent transformer models and sometimes even outperform them, as in the case of for global diversity. This shows that unconditional LLMs avoid repetitions at a local level whereas, at the sample level, they generate poems that are lexically much more similar to each other than poems within the human sample. ††todo: explain the differences between MATTR and MTLD
Semantic Similarity
Model | Within (%) | Across (%) |
HUMAN | 55.0 / 48.2 | - |
DeepSpeare | 59.5 / 52.2 | 67.8 / 60.8 |
SA | 55.8 / 49.6 | 65.9 / 59.4 |
ByGPT5 | 58.4 / 53.2 | 68.1 / 61.5 |
ByGPT5 | 58.2 / 52.7 | 67.9 / 61.6 |
GPT2 | 64.5 / 59.5 | 69.3 / 63.9 |
GPT2 | 63.6 / 57.6 | 70.1 / 63.3 |
GPTNeo | - / 62.2 | - / 63.8 |
GPTNeo | - / 60.9 | - / 63.9 |
LLaMA2 | 61.0 / 59.4 | 68.5 / 64.2 |
LLaMA2 | 62.3 / 58.0 | 68.9 / 62.9 |
LLaMA2 | 61.2 / 58.4 | 69.1 / 63.8 |
ByGPT5 | 58.4 / 52.2 | 67.7 / 60.8 |
ByGPT5 | 57.9 / 50.9 | 67.6 / 60.3 |
GPT2 | 64.3 / 59.2 | 70.1 / 64.3 |
GPT2 | 62.6 / 57.4 | 69.7 / 63.1 |
GPTNeo | - / 58.9 | - / 64.0 |
GPTNeo | - / 60.3 | - / 62.9 |
LLaMA2 | 66.9 / 57.3 | 69.3 / 64.0 |
LLaMA2 | 63.3 / 58.5 | 69.5 / 62.9 |
LLaMA3 | 59.6 / 58.2 | 68 / 62.3 |
Table 3 presents results for the semantic (cosine) similarity of quatrains: (i) within human and model-generated samples, and (ii) across generated samples and the human data. These results generally confirm the trends we found for global lexical diversity discussed above. None of the models generates a sample of poems with a within-sample diversity as low as the human with-sample diversity. SA is the model that achieves the lowest within-sample similarity and the lowest across-sample similarity, suggesting that it deviates most from the patterns in the human training data. Note that SA also achieved the best results in global lexical diversity (MATTR and MTLD in Table 2).††todo: SZ: I am not entirely sure about the across-sample similarities: is it necessarily a good thing if the models have low similarity with the human sample?. Moreover, the results on semantic similarity confirm the trends we observed with model size for lexical diversity, but disconfirm the trends for the effect of conditioning. Thus, we do not see a consistent trend for conditioned models generating samples with lower similarity/higher diversity. For model size, on the other hand, we observe a general trend towards larger models outperforming their smaller counterparts.
7 Discussion
Sampling/Searching
We deploy various decoding strategies to determine to what extent these can alter the various aspects of diversity in the generated poems. We use different combinations of temperature-based sampling (ackley1985learning), Nucleus sampling (Top-p) (holtzman2019curious) and Top-k sampling (fan-etal-2018-hierarchical) as sampling strategies and further deploy two variants of contrastive search (su2022contrastive). Results indicate that the various techniques can only slightly increase diversity in one or more aspects. Moreover, aggressive sampling often leads to output degeneration, causing the models to (partially) repeat verses in a quatrain. We provide some examples in the appendix, see Tables 7 to 12 as well as Figures 10 to 15.
Which is the most diverse model?
We have seen that unconditioned LLMs exhibit poor results with regard to different dimensions of diversity: they do not rhyme, are lexically underdiverse and do not show sufficient semantic variation. However, character-level models are more diverse than word level models. Style-conditioned models perform better regarding rhyming, semantic variation, and lexical variation but word level style-conditioned models are prone to severe memorization from the training data, in particular when the model is large and the training set is small. Character-level style-conditioned LLMs produce overall best diversity results and do not deteriorate as a function of model/training data size. In terms of diversity, poetry-specific Deepspeare performs similar as character-level LLMs but requires more modeling effort from human experts (e.g., in developing rhyming components).
8 Conclusion
To date, evaluation of automatic poetry generation has almost exclusively focused on human evaluation and forms of the Turing test. Our work shows that an automatic assessment of the diversity of generated poems covers an important blind spot of existing studies. Our evaluations shed light on the fact that none of the state-of-the-art poetry generators is able to match the level of diversity in human poems, confirming previous evaluations of diversity in other NLG tasks ippolito-etal-2019-comparison; schuz-etal-2021-diversity; stasaski-hearst-2022-semantic. Our study also adds a new dimensions to previous work on diversity, by showing that diversity on the level of rhyming is particularly hard to achieve for neural generators and interacts with other dimensions of diversity in poetry generation, i.e., style conditioned LLMs do not only achieve a better match with human rhyme and length distributions, but also higher lexical and semantic diversity . We also find that memorization — a general and widely discussed limitation of LLMs (carlini2021extracting) — is a potential issue in poetry generation, especially for certain combinations of model sizes and finetuning schemes, complementing existing studies in this area mireshghallah-etal-2022-empirical.
We release all code upon acceptance.
Limitations
Our work evaluates a range of existing state-of-the-art approaches, such as poetry-specific models like Deepspeare or pretrained LLMs. These models differ in various ways, with respect to their architecture, training scheme, pretraining, and the type of data they expect during training and/or finetuning. In light of these differences, it is difficult to isolate exactly how different aspects of a poetry generator impact on the diversity of its outputs. While our work investigated the influence of the model architecture on a high level (character vs. word), further aspects — and in particular pre-training — may be worth investigating in future work.
Generally, our work is concerned with the evaluation of NLG systems; evaluation methods and evaluation metrics (zhao-etal-2019-moverscore; Zhang2020BERTScore; NEURIPS2021_e4d2b6e6; chen_menli:2023; peyrard-etal-2021-better) are a well-known and notorious issue in this research field. While a lot of recent work has aimed at improving common practices in human evaluation (belz-etal-2023-missing) or advancing the study of metrics for quality or fluency of NLG outputs, the evaluation of diversity is comparatively under-researched. In this work, we aimed at providing a range of metrics assessing different aspects of diversity, but could not cover all potentially interesting ways of measuring diversity. Here, future work could look at further aspects of formal and structural diversity (e.g. at the level of syntax, or meter), or other aspects of semantic diversity (e.g. topical diversity, rhetorical figures). Future work could also consider more (diverse) languages and other genres and datasets for poetry.
Ethics Statement
Often, the discussion of creative AI systems in public discourse is surrounded by misconceptions, hypes and even myths (veale2012exploding). Our work contributes to a careful operationalization and objective assessment of the creative capbalities of AI systems in the area of poetry generation.
All the datasets, models and code used in this work are publicly available or will be made available upon publication. We have not collected private or sensitive data and have only used language models with free access, such that our experiments can be fully replicated by anyone.
Appendix A Appendix
Model | EN | DE |
GPT2-small | 124M | 124M |
GPT2-large | 774M | 774M |
GPTneo-small | 125M | - |
GPTneo-xl | 1.3B | - |
byGPT5-base | 140M | 140M |
byGPT5-medium | 290M | 290M |
model | |||||||
de | HUMAN | 1.00 | 4 | 65 | 24.40 | 23 | 6.39 |
de | DeepSpeare | 0.63 | 14 | 30 | 21.69 | 22 | 2.45 |
de | SA | 0.88 | 10 | 44 | 24.44 | 24 | 5.36 |
de | ByGPT5 | 0.84 | 9 | 43 | 22.11 | 22 | 4.86 |
de | ByGPT5 | 0.79 | 9 | 40 | 21.09 | 21 | 4.59 |
de | GPT2 | 0.59 | 9 | 32 | 19.18 | 19 | 3.54 |
de | GPT2 | 0.73 | 13 | 41 | 21.98 | 22 | 3.55 |
de | LLaMA2 | 0.57 | 9 | 31 | 18.84 | 19 | 3.29 |
de | LLaMA2 | 0.55 | 9 | 30 | 18.73 | 19 | 3.17 |
de | LLaMA3 | 0.74 | 12 | 40 | 21.39 | 21 | 3.99 |
de | ByGPT5 | 0.82 | 11 | 47 | 22.38 | 22 | 4.98 |
de | ByGPT5 | 0.81 | 9 | 45 | 21.78 | 21 | 5.17 |
de | GPT2 | 0.70 | 11 | 37 | 20.68 | 20 | 3.56 |
de | GPT2 | 0.79 | 14 | 45 | 24.14 | 24 | 4.38 |
de | LLaMA2 | 0.83 | 12 | 49 | 24.22 | 23 | 5.41 |
de | LLaMA2 | 0.62 | 12 | 34 | 20.18 | 20 | 2.84 |
de | LLaMA3 | 0.76 | 10 | 47 | 21.69 | 21 | 4.14 |
en | HUMAN | 1.00 | 4 | 67 | 28.06 | 28 | 6.26 |
en | DeepSpeare | 0.57 | 15 | 33 | 23.85 | 24 | 2.85 |
en | SA | 0.92 | 12 | 52 | 27.36 | 27 | 5.38 |
en | ByGPT5 | 0.80 | 12 | 44 | 25.30 | 25 | 5.09 |
en | ByGPT5 | 0.77 | 11 | 47 | 24.97 | 25 | 4.87 |
en | GPT2 | 0.69 | 13 | 55 | 24.11 | 24 | 4.48 |
en | GPT2 | 0.72 | 13 | 56 | 24.74 | 24 | 4.94 |
en | GPTNeo | 0.55 | 11 | 55 | 22.67 | 22 | 3.89 |
en | GPTNeo | 0.48 | 13 | 34 | 21.93 | 22 | 3.16 |
en | LLaMA2 | 0.87 | 15 | 75 | 28.60 | 27 | 7.52 |
en | LLaMA2 | 0.67 | 12 | 54 | 23.95 | 24 | 4.50 |
en | LLaMA3 | 0.59 | 14 | 60 | 23.20 | 23 | 4.23 |
en | ByGPT5 | 0.85 | 13 | 42 | 26.21 | 26 | 4.96 |
en | ByGPT5 | 0.84 | 14 | 42 | 25.85 | 25 | 4.84 |
en | GPT2 | 0.86 | 17 | 61 | 28.37 | 27 | 6.18 |
en | GPT2 | 0.83 | 16 | 70 | 27.82 | 27 | 6.15 |
en | GPTNeo | 0.74 | 16 | 49 | 25.13 | 24 | 4.47 |
en | GPTNeo | 0.53 | 12 | 35 | 22.26 | 22 | 3.36 |
en | LLaMA2 | 0.70 | 17 | 74 | 33.55 | 32 | 7.83 |
en | LLaMA2 | 0.81 | 15 | 56 | 26.92 | 26 | 5.80 |
en | LLaMA3 | 0.78 | 16 | 65 | 27.12 | 26 | 5.35 |
A.1 Rhyme
DE | EN | |||||
Model | Entropy | KL Divergence | Cross Entropy | Entropy | KL Divergence | Cross Entropy |
HUMAN | 2.90 | 0.00 | 2.90 | 3.10 | 0.00 | 3.10 |
DeepSpeare | 2.97 | 0.55 | 3.52 | 3.16 | 0.48 | 3.64 |
SA | 3.14 | 1.43 | 4.57 | 3.22 | 1.17 | 4.39 |
ByGPT5 | 2.89 | 1.23 | 4.12 | 2.92 | 1.08 | 4.00 |
ByGPT5 | 3.13 | 1.09 | 4.21 | 2.91 | 1.13 | 4.04 |
GPT2 | 2.86 | 1.26 | 4.13 | 2.97 | 1.06 | 4.03 |
GPT2 | 3.16 | 1.13 | 4.29 | 2.99 | 1.03 | 4.03 |
GPTNeo | - | - | - | 2.80 | 1.18 | 3.99 |
GPTNeo | - | - | - | 3.16 | 0.96 | 4.11 |
LLaMA2 | 2.93 | 1.18 | 4.11 | 3.24 | 0.71 | 3.95 |
LLaMA2 | 3.18 | 1.04 | 4.21 | 3.24 | 0.71 | 3.95 |
LLaMA3 | 3.27 | 0.83 | 4.10 | 3.45 | 0.56 | 4.02 |
ByGPT5 | 3.17 | 0.67 | 3.84 | 3.22 | 0.83 | 4.05 |
ByGPT5 | 3.16 | 0.58 | 3.75 | 3.38 | 0.54 | 3.92 |
GPT2 | 2.98 | 0.99 | 3.97 | 3.41 | 0.61 | 4.02 |
GPT2 | 3.11 | 1.04 | 4.14 | 3.22 | 0.85 | 4.07 |
GPTNeo | - | - | - | 3.43 | 0.45 | 3.89 |
GPTNeo | - | - | - | 3.29 | 0.83 | 4.11 |
LLaMA2 | 2.69 | 1.33 | 4.02 | 2.89 | 0.95 | 3.85 |
LLaMA2 | 3.11 | 0.71 | 3.83 | 2.67 | 1.07 | 3.74 |
LLaMA3 | 2.98 | 1.06 | 4.04 | 2.58 | 0.94 | 3.52 |
Appendix B Decoding methods
Method | GPT2-large | p-GPTneo-xl | p-byGPT5-medium |
0.002 | 0.89 | 0.006 | |
0.000 | 0.96 | 0.006 | |
0.000 | 0.98 | 0.004 | |
0.002 | 0.99 | 0.010 | |
0.004 | 0.98 | 0.004 | |
0.012 | 0.99 | 0.020 | |
0.002 | 0.99 | 0.010 | |
0.000 | 0.99 | 0.008 | |
0.002 | 0.98 | 0.004 | |
0.004 | 0.99 | 0.012 | |
0.004 | 0.94 | 0.004 |
Method | mean | std | min | max | |
0.002 | 0.55 | 0.06 | 0.33 | 0.77 | |
0.000 | 0.57 | 0.06 | 0.37 | 0.82 | |
0.000 | 0.58 | 0.06 | 0.39 | 0.72 | |
0.002 | 0.57 | 0.07 | 0.38 | 0.81 | |
0.004 | 0.57 | 0.06 | 0.37 | 0.97 | |
0.012 | 0.59 | 0.07 | 0.44 | 0.99 | |
0.002 | 0.56 | 0.06 | 0.37 | 0.91 | |
0.000 | 0.58 | 0.07 | 0.33 | 0.84 | |
0.002 | 0.57 | 0.07 | 0.42 | 0.95 | |
0.004 | 0.57 | 0.07 | 0.39 | 0.98 | |
0.004 | 0.56 | 0.06 | 0.34 | 0.94 |
Method | mean | std | min | max | h | |
29.83 | 13.64 | 4 | 150 | 0.75 | 5.46 | |
37.77 | 20.58 | 4 | 150 | 0.87 | 4.48 | |
35.67 | 8.85 | 13 | 64 | 0.86 | 1.89 |
Model | lang | real rhymes | repetitions |
en | 0.72 | 0.00 | |
en | 0.84 | 0.00 | |
en | 0.44 | 0.08 | |
en | 0.53 | 0.21 | |
de | 0.57 | 0.00 | |
de | 0.66 | 0.00 |
method | real rhymes | repetitions |
0.156 | 0.022 | |
0.196 | 0.052 | |
0.164 | 0.071 | |
0.176 | 0.220 | |
0.170 | 0.103 | |
0.181 | 0.124 | |
0.146 | 0.030 | |
0.168 | 0.131 | |
0.191 | 0.112 | |
0.175 | 0.153 | |
0.194 | 0.087 |
Method | ATTR | MATTR | MTLD |
0.804 | 0.794 | 64.679 | |
0.757 | 0.744 | 44.942 | |
0.716 | 0.717 | 36.561 | |
0.537 | 0.537 | 17.297 | |
0.720 | 0.690 | 33.151 | |
0.419 | 0.461 | 14.730 | |
0.586 | 0.622 | 21.757 | |
0.611 | 0.588 | 21.005 | |
0.766 | 0.754 | 46.107 | |
0.642 | 0.637 | 24.289 | |
0.797 | 0.779 | 58.490 |

### v2

@@
 Creative NLG, such as automatic poetry generation, is a fascinating niche in this area.
-While most previous research has focused on forms of the Turing test when evaluating automatic poetry generation — can humans distinguish between automatic and human generated poetry — we evaluate the diversity of automatically generated poetry, by comparing distributions of generated poetry to distributions of human poetry along structural, lexical, semantic and stylistic dimensions, assessing different model types (word vs.
+While most previous research has focused on forms of the Turing test when evaluating automatic poetry generation — can humans distinguish between automatic and human generated poetry — we evaluate the diversity of automatically generated poetry (with a focus on quatrains), by comparing distributions of generated poetry to distributions of human poetry along structural, lexical, semantic and stylistic dimensions, assessing different model types (word vs.
 character-level, general purpose LLMs vs.
-poetry-specific models) and types of fine-tuning (conditioned vs.
+poetry-specific models), including the very recent LLaMA3-8B, and types of fine-tuning (conditioned vs.
 unconditioned).
-We find that current automatic poetry systems are considerably underdiverse along all dimensions — they tend to memorize, do not rhyme sufficiently, are semantically too uniform and even do not match the length distribution of human poetry.
-Among all models explored, character-level style-conditioned models perform slightly better.
-Our identified limitations may serve as the basis for more genuinely creative future poetry generation models.
+We find that current automatic poetry systems are considerably underdiverse along multiple dimensions — they often do not rhyme sufficiently, are semantically too uniform and even do not match the length distribution of human poetry.
+Our experiments reveal, however, that style-conditioning and character-level modeling clearly increases diversity across virtually all dimensions we explore.
+Our identified limitations may serve as the basis for more genuinely diverse future poetry generation models.111Code + data: https://github.com/hgroener/diversity_in_poetry_generation
 Evaluating Diversity in Automatic Poetry Generation
-Anonymous ACL submission
+Yanran Chen1, Hannes Gröner2, Sina Zarrieß2, Steffen Eger1 1 NLLG, University of Mannheim & University of Technology Nuremberg (UTN); https://nl2g.github.io/ 2 Computational Linguistics, Bielefeld University yanran.chen@uni-mannheim.de {hannes.groener,sina.zarriess}@uni-bielefeld.de steffen.eger@utn.de
 1 Introduction
 A key aspect of creative language generation is the ability to create new, original and interesting text, cf.
-colton2012computational; gatt2018survey; yi2020mixpoet; Elgammal2017CANCA.
-To date, extremely little attention has been given to the evaluation of originality and creativity in recent creative text generation models such as those for automatic poetry generation, despite renewed interest in the context of recent LLMs franceschelli2023creativity.
-In fact, existing automatic poetry generation models are typically not evaluated regarding how different generated poems are from existing poems in the training set but with the Turing test: can humans distinguish whether a poem is human authored or automatically generated (hopkins-kiela-2017-automatically; lau2018deep; Manjavacas2019ARS)?
+Colton et al.
+(2012); Gatt and Krahmer (2018); Yi et al.
+(2020); Elgammal et al.
+(2017).
+To date, extremely little attention has been given to the evaluation of originality and creativity in recent creative text generation models such as those for automatic poetry generation, despite renewed interest in the context of recent LLMs Franceschelli and Musolesi (2023).
+In fact, existing automatic poetry generation models are typically not evaluated regarding how different generated poems are from existing poems in the training set but with the Turing test: can humans distinguish whether a poem is human authored or automatically generated (Hopkins and Kiela, 2017; Lau et al., 2018; Manjavacas et al., 2019)?
 However, this form of Turing test and other similar forms of human evaluation may contain an overlooked risk of failure: namely, if the automatically generated instances are (near-)copies of training data instances.
-In this work, we fill this gap and evaluate, for the first time, automatic poetry generation systems for their diversity.
-As human evaluation is generally not well suited to assess diversity hashimoto-etal-2019-unifying, we automatically measure diversity by comparing distributions of generated and existing poems along formal, semantic and stylistic dimensions.
+In this work, we fill this gap and evaluate, for the first time, (fine-tuned) automatic poetry generation systems in terms of their diversity.
+As human evaluation is generally not well suited to assess diversity Hashimoto et al.
+(2019), we automatically measure diversity by comparing distributions of generated and existing poems along formal, semantic and stylistic dimensions.
 This yields much better evidence of the models’ creative capabilities in contrast to being mere ‘stochastic parrots’.
-Our main contributions are: (i) we conceptualize diversity of poetry generation systems along different dimensions: diversity on the structural, lexical, semantic and stylistic level; (ii) we assess different types of automatic poetry generation systems for diversity: general purpose word and character-level LLMs, both unconditioned and style-conditioned ones, on the one hand, and poetry-specific models, on the other hand; (iii) we evaluate each class of model for diversity across the different dimensions, by comparing the distribution of the human authored training data set to the distribution of generated poems.
+Our main contributions are: (i) we conceptualize diversity of poetry generation systems along different dimensions: diversity on the structural (e.g., length), stylistic (e.g., rhyming), lexical and semantic level; (ii) we assess different types of automatic poetry generation systems for diversity: general purpose word- and character-level LLMs, both unconditioned and style-conditioned ones, on the one hand, and poetry-specific models, on the other hand; (iii) we evaluate each class of model for diversity across the different dimensions, by comparing the distribution of the human authored training data set to the distribution of generated poems.
 We find that on a distributional level, generated poems are considerably different from human ones.
-Concerning general purpose LLMs, some of them exhibit very high risk of memorization — an extreme form of lack of diversity — and this depends on the size of the training data set, the size and type of the LLM, and the type of training, as we show.
 Character-level style-conditioned general-purpose LLMs are most diverse.
-Our work prepares the groundwork for truly creative generative AI models (veale2020leaps) and also has implications for the detection of generative AI (Sadasivan2023CanAT).
 2 Related Work
@@
 Diversity
-Building systems able to generate diverse output has been a long-standing concern in NLG research (reiter-sripada-2002-squibs; van-deemter-etal-2005-squibs; foster-white-2007-avoiding) and remains a central issue in neural NLG (holtzman2019curious).
-The need for careful analysis of NLG systems diversity – beyond an assessment of the quality or fluency of single-best generation outputs – has been widely acknowledged (gatt2018survey; hashimoto-etal-2019-unifying; mahamood-zembrzuski-2019-hotel; celikyilmaz2020evaluation; tevet-berant-2021-evaluating; schuz-etal-2021-diversity).
-A well-known finding from this line of research is that neural NLG systems typically face a quality-diversity trade-off (ippolito-etal-2019-comparison; caccia2020; wiher2022decoding): their outputs are either well-formed and fluent or diverse and variable.
-Work on evaluating diversity of NLG typically uses automatic metrics that quantify to what extent different outputs by the same system vary (hashimoto-etal-2019-unifying).
-In practice, though, evaluations of diversity in NLG differ widely across tasks (tevet-berant-2021-evaluating) and even adopt different notions of diversity (zarriess-etal-2021-decoding).
+Building systems able to generate diverse output has been a long-standing concern in NLG research (Reiter and Sripada, 2002; van Deemter et al., 2005; Foster and White, 2007) and remains a central issue in neural NLG (Holtzman et al., 2019).
+The need for careful analysis of NLG systems’ diversity – beyond an assessment of the quality or fluency of single-best generation outputs – has been widely acknowledged (Gatt and Krahmer, 2018; Hashimoto et al., 2019; Mahamood and Zembrzuski, 2019; Celikyilmaz et al., 2020; Tevet and Berant, 2021; Schüz et al., 2021).
+A well-known finding from this line of research is that neural NLG systems typically face a quality-diversity trade-off (Ippolito et al., 2019; Caccia et al., 2020; Wiher et al., 2022): their outputs are either well-formed and fluent or diverse and variable.
+Work on evaluating diversity of NLG typically uses automatic metrics that quantify to what extent different outputs by the same system vary (Hashimoto et al., 2019).
+In practice, though, evaluations of diversity in NLG differ widely across tasks (Tevet and Berant, 2021) and even adopt different notions of diversity (Zarrieß et al., 2021).
 At the same time, most of these notions focus on lexical or semantic aspects of diversity, e.g., local lexical diversity.
-For instance, ippolito-etal-2019-comparison compare decoding methods in dialog generation and image captioning, assessing lexical overlaps in -best NLG outputs for the same input.
+For instance, Ippolito et al.
+(2019) compare decoding methods in dialogue generation and image captioning, assessing lexical overlaps in -best NLG outputs for the same input.
+Chakrabarty et al.
+(2022) simply measure the local lexical diversity in automatic generated poems in terms of distinct unigrams.
 Global lexical diversity, on the other hand, measures whether the NLG system generates different outputs for different inputs.
-For instance, van-miltenburg-etal-2018-measuring define the global diversity of image captioning systems as their ability to generate different captions for a set of inputs, using metrics like the number of types in the output vocabulary, type-token ratio, and the percentage of novel descriptions.
-Similarly, hashimoto-etal-2019-unifying view diversity as related to the model’s ability to generalize beyond the training set, i.e., generate novel sentences.
+For instance, van Miltenburg et al.
+(2018) define the global diversity of image captioning systems as their ability to generate different captions for a set of inputs, using metrics like the number of types in the output vocabulary, type-token ratio, and the percentage of novel descriptions.
+Similarly, Hashimoto et al.
+(2019) view diversity as related to the model’s ability to generalize beyond the training set, i.e., generate novel sentences.
 Besides lexical diversity, work on open-ended or creative text generation tasks has been interested in diversity at a more general semantic level.
-For instance, zhang2018 and stasaski-hearst-2022-semantic aim at building dialogue systems that generate entertaining and semantically diverse responses in chit-chat dialog, where the goal is to avoid “safe and bland” responses that “average out” the sentences observed in the training set.
-Here, semantic diversity has been measured, e.g., with the help of embedding-based similarity (du-black-2019-boosting).
-In our work on diversity in poetry generation, we complement these lexical and semantic aspects of diversity with aspects of formal diversity.
+For instance, Zhang et al.
+(2018) and Stasaski and Hearst (2022) aim at building dialogue systems that generate entertaining and semantically diverse responses in chit-chat dialog.
+Here, semantic diversity has been measured, e.g., with the help of embedding-based similarity (Du and Black, 2019).
+In our work on diversity in poetry generation, we complement both lexical and semantic aspects of diversity with aspects of formal diversity.
 We thus explore whether automatic poetry generation systems are able to capture the ‘full bandwidth’ of realizations of poetry found in the data distribution with which they have been trained, focusing mostly on global diversity.
@@
 Automatic poetry generation is a long standing dream of AI research, dating back at least to the mid 20th century (e.g., Theo Lutz’ Stochastische Texte).
-While early modern systems were heavily hand-engineered gervas2001expert, more recent approaches are all trained on collections of human poetry lau2018deep; jhamtani2019learning; agarwal-kann-2020-acrostic but still extensively utilize human guidance e.g.
-to enforce formal characteristics of poetry such as rhyming wockener-etal-2021-end.
-belouadi2022bygpt5 have recently released a character-level decoder-only LLM (ByGPT5) capable of learning style-constraints such as rhyming without human involvement in model design.
+While early modern systems were heavily hand-engineered Gervás (2001), more recent approaches are all trained on collections of human poetry Lau et al.
+(2018); Jhamtani et al.
+(2019); Agarwal and Kann (2020) but still extensively utilize human guidance e.g.
+to enforce formal characteristics of poetry such as rhyming Wöckener et al.
+(2021).
+Belouadi and Eger (2023) have recently released a character-level decoder-only LLM (ByGPT5) capable of learning style-constraints such as rhyming without human involvement in model design.
+Chakrabarty et al.
+(2022) propose a collaborative system for poetry, which can follow human instructions to write poems.
+They measure creativity of the generated poems via crowd workers, who decide which of two poems is more creative.
+While Chakrabarty et al.
+(2022) do not define creativity, it could be considered as generating novel poems outside the training data set; in contrast, we measure diversity by assessing whether poetry generation systems generate outputs that are as diverse as their human training data.
 In our work, we explore varying poetry generation models with regard to diversity: poetry-specific models that use hand-engineered architectures as well as general purpose LLMs, including ByGPT5.
@@
 We first conceptualize diversity in poetry generation using formal and semantic criteria.
-As our dataset, we use QuaTrain (belouadi2022bygpt5) consisting of quatrains (in English and German).
-We describe it in more detail in §5.1 below.
 Memorization.
-In poetry, as in other forms of art, creativity (sternberg1999handbook) plays a central role.
+In poetry, as in other forms of art, creativity (Sternberg, 1999) plays a central role.
 A basic aspect of creativity is the models’ ability to generate poems that are different from the training data, i.e.
 have not been memorized as a whole.
-Therefore, we consider a low or minimal degree of memorization as a pre-requisite for diversity and analyze the portion of generated poems that are (near-)copies from the training data.
-To examine memorization, we proceed as in belouadi2022bygpt5.
-We apply the Ratcliff-Obershelp similarity (ratcliff1988pattern) to compare each poem in a sample with poems in the training corpus.
+To examine memorization, we proceed as in Belouadi and Eger (2023).
+We apply the Ratcliff-Obershelp similarity (Ratcliff et al., 1988) to compare each poem in a sample with poems in the training corpus.
 If a generated quatrain exhibits a similarity score of 0.7 with a quatrain in the training data, we classify it as memorized.
@@
 Higher thresholds for classification are used for these finer-grained comparison levels, as shorter texts have higher chances of being more similar in general.
-11todo: 1YC: maybe add a table in the appendix to demonstrate the couplets/verses classified as memorized with threshold 0.7.
 Specifically, a verse with a similarity score 0.9 or a couplet 0.8 is considered as memorized.
 We define the memorization score of a sample as the proportion of memorized quatrains in that sample.
-How much LLMs memorize from their training data has been a question of central concern recently mccoy-etal-2023-much.
+How much LLMs memorize from their training data has been a question of central concern recently McCoy et al.
+(2023).
 Poem length.
@@
 We report mean length, standard deviation, minimal and maximal length of samples.
-We additionally deploy distance measures between training data distribution and generated samples, in particular, a metric called histogram intersection (swain1991color), which measures the intersection area of two normalized histograms (and therefore returns values between 0 and 1).
+We additionally deploy distance measures between training data distribution and generated samples, in particular, a metric called histogram intersection (Swain and Ballard, 1991), which measures the intersection area of two normalized histograms (and therefore returns values between 0 and 1).
 Rhyme patterns.
@@
 Diversity can then be assessed by comparing rhyme distributions between generated samples and training data.
-In order to classify rhymes in our samples, we use the same classifier used to annotate QuaTrain.
+In order to classify rhymes in our samples, we use the same classifier used to annotate QuaTrain (Belouadi and Eger, 2023).
 We distinguish between true rhymes, which involve different words, and repetitions, which refer to rhymes based on the same word.
@@
 We use the following metrics to measure the lexical diversity for both the training data and the generated samples: (i) Averaged type token ratio (ATTR).
-We calculate ATTR as the average of all type token ratios (richards1987type) (TTRs) for each quatrain in a sample, i.e.
+We calculate ATTR as the average of all type token ratios (Richards, 1987) (TTRs) for each quatrain in a sample, i.e.
 as a measure of local lexical diversity.
 (ii) Moving average type token ratio (MATTR).
-The MATTR (covington2010cutting) acts on the corpus level and calculates a moving average by sliding through the corpus using a window of fixed size.
+The MATTR (Covington and McFall, 2010) acts on the corpus level and calculates a moving average by sliding through the corpus using a window of fixed size.
 We deploy this metric as a measure of global lexical diversity.
 (iii) Measure of textual, lexical diversity (MTLD).
-The MTLD (mccarthy2005assessment) is calculated as the average length of a substring that maintains a specified TTR level.
+The MTLD (McCarthy, 2005) is calculated as the average length of a substring that maintains a specified TTR level.
 MTLD is deployed to measure lexical diversity on a global scale.
@@
 Even if a poetry generation system does not directly copy data from the training data, the generated poems may still be semantically very similar to the training data distribution.
-We employ a multilingual distilled version of Sentence-BERT (SBERT) reimers-gurevych-2019-sentence as dense vector representations to measure semantic similarity between poems: (i) across the human train set and the generated poems, (ii) within human and generated poems.
+We employ a multilingual distilled version of Sentence-BERT (SBERT) Reimers and Gurevych (2019) as dense vector representations to measure semantic similarity between poems: (i) across the human train set and the generated poems, (ii) within human and generated poems.
 In particular, for each generated quatrain, we note down the similarity value of the most similar human quatrain, then report the average over all those maximum similarity values.
 We proceed analogously within the human training data and within the automatically generated poems.
-4 Models
-Our experiments use 2 different model classes.
-4.1 Poetry-specific models
-Deepspeare.
-Deepspeare lau2018deep is specifically designed for poetry generation.
-Its core architecture consists of an LSTM language model, a pentameter model (specifically designed to learn iambic meter) and a rhyme model.
-During training, it takes sonnets as input data (three quatrains followed by a couplet) but ultimately processes the contained quatrains by splitting any given sonnet.
-The rhyme model processes ending words of quatrain verses and uses a margin-based loss to discriminate between rhyming and non-rhyming words.
-It is not limited to specific rhyme patterns but assumes that rhymes exist in the data.
-At inference time, Deepspeare generates quatrains.
-Structured Adversary.
-Like Deepspeare, Structured Adversary (SA) jhamtani2019learning incorporates different components: an LSTM language model and a discriminator used to decide whether line endings are typical for poetry.
-Both components are organized in an adversarial setup, where the language model acts as a generator, trying to generate poems that are misclassified by the discriminator, while the discriminator is trained to distinguish generated poems from real ones.
-SA is trained with sonnets as input data.
-At inference time, it generates quatrains.
-4.2 General purpose LLMs
-All models in this category are decoder-only transformer architectures. In our experiments, we train them in an unconditioned and style-conditioned manner (see Section 5.2).
-GPT2
-GPT2 radford2019language is the last GPT model made publicly available.
-It is a large word level transformer-based language model pretrained on approximately 40 GB of text.
-Four different model versions were released, with the number of parameters ranging from 125 million to 1.5 billion for the largest.
-In this work, we deploy two model versions: GPT2-small (125M parameters) and GPT2-large (774M parameters) for both English and German.
-GPTneo
-GPTneo (black-etal-2022-gpt) is an open-source token level LLM by EleutherAI (https://www.eleuther.ai/) with the aim to provide publicly available replications of GPT3.
-It is pre-trained on 825 GB of text data.
-Currently, four versions have been released, with the number of parameters ranging from 125 million up to 20 billion.
-We deploy GPTneo-small and GPTneo-xl with 125M and 1.3B parameters for English.
-GPTneo is not available for German.
-LLaMA2/3
-ByGPT5
-ByGPT5 (belouadi2022bygpt5) is a decoder-only character level LLM based on the encoder-decoder character level model byT5 (xue-etal-2022-byt5) where the encoder part of byT5 is completely removed, reducing the number of parameters by 75%.
-The remaining decoder-only model is then pretrained using OpenWebText for English (38GB text data) and CC100 (conneau-etal-2020-unsupervised) (67GB text data) for German.
-Three versions are released for both English and German, with model sizes ranging from 73 to 298M parameters.
-We use ByGPT5-base (140M params) and ByGPT5-medium (290M) for both English and German.
-5 Experimental Setup
-5.1 Training Data
+4 Experiment Setup
+Data
 | DE | EN | |||
@@
 | Total | 282,048 | 80,584 | 201,856 | 57,672 |
-We use QuaTrain, a large dataset of quatrains published by belouadi2022bygpt5.
-It consists of English and German quatrains and has been generated by aggregating different publicly available poetry datasets.
-QuaTrain contains human written quatrains but mixes them synthetically: every sequence of four consecutive lines from the underlying human data are included in order to increase dataset size.
-QuaTrain is automatically annotated for meter and rhyme using high-quality classifers (especially for rhyme).
-22todo: 2YC: @Hannes, could you please add here how you preprocessed the data to drop the redundant ones?
-H: on it!
+We use the QuaTrain dataset published by Belouadi and Eger (2023), which consists of English and German quatrains from different publicly available poetry datasets.
+The dataset contains human written quatrains but mixes them synthetically: every sequence of four consecutive lines from the underlying human data are included in order to increase dataset size.
+Besides, it is automatically annotated for meter and rhyme using high-quality classifers (especially for rhyme).
 Because our focus lies on the diversity of model outputs, we have to avoid repetitions in the training data created by the data augmentation methods used in its creation.
 To avoid lines appearing multiple times, we first parse the dataset sequentially, eliminating quatrains that overlap the preceding one.
-Because this method does not eliminate all overlaps, we then use an heuristic, deleting the ten percent of the quatrains which have the biggest overlap with other quatrains until there is no overlap remaining.
-We use a ratio of 9:1 to split the data into train and dev sets; we do not keep a test set since no held-out human data is needed for generation or evaluation.
-Further, as DeepSpeare and SA are designed to process sonnets and/or limerick data, we create pseudo sonnets for them using the corresponding data splits.
-Specifically, for each sonnet, we randomly draw three quatrains and one couplet from the QuaTrain datasets, ensuring that each comes from a different original quatrain.
-Table 1 provides the resulting data statistics.
-5.2 Training
-DeepSpeare.
-DeepSpeare (lau2018deep) leverages pretrained static word vectors.
-We use QuaTrain and SonNet to train our own Word2vec embeddings (mikolov2013efficient) and the final sonnet models respectively.
-For the sonnet model training, we use a batch size of 128 and apply early stopping with a patience of 5 epochs; default settings are maintained for the other hyperparameters.
-SA.
-We use the same word vectors and training data splits as for DeepSpeare.
-Training SA involves 1) pretraining the discriminator’s encoder using a publicly available pronouncing dictionary ; 2) training the LM component; 3) training a final aggregated model in a generative adversarial setup.
-We train the discriminators with a batch size of 128, the LMs with a batch size of 64, and the final sonnet models with a batch size of 128; here, we also implement early stopping with a patience of 5 epochs.
-Unconditioned LLMs.
-In this setup, we fine-tune our decoder-only LLMs in an unconditioned manner: we process quatrains during training without passing any information about rhyme (or meter).
-All models except GPTneo (being available only in English) are trained both in English and German.
-We fine-tune all models (English and German) for 50 epochs, with early stopping using a patience of 3 epochs.
-Style-conditioned LLMs.
-In contrast to unconditioned training, we provide information about rhyme (and meter) by prepending special style tokens to each quatrain during training.
-This follows the setup of belouadi2022bygpt5 and makes models explicitly aware of different rhyme schemes.
-As for the unconditioned variants, all models except GPTneo are trained in English and German.
-We use the same hyperparameter settings as for the unconditioned trained LLMs.
-Summary.
-We end up with 36 models that can be assigned to three categories: 1) Poetry specific LSTM-based models (DeepSpeare and SA).
-Besides a language model part, these models incorporate additional specialized components to handle poetry-specific stylistic features such as rhyme.
-We have 2 models for each language.
-2) Unconditioned LLMs (transformer-based decoder-only general purpose LLMs).
-These models do not possess any specialized architecture for poetry.
-No information about meter or rhyme has actively been passed during training.
-We have two subcategories: word and character level models.
-The first group (the GPT and LLaMA series) processes data on the word/subword level.
-ByGPT5 represents the character-level group.
-We have 9 models for English and 7 for German.
-3) Style-conditioned LLMs.
-These have the same architecture, models, and subgroups as category 2.
-Information about rhyme (and meter) is passed in the form of special tokens during training (only).
-We use the following notation to distinguish the size and training manners of the LLMs: a subscript of S/L indicates whether it is a smaller/larger version, and a superscript of “con” stands for conditioned training.
+Because this method does not eliminate all overlaps, we then use a heuristic, deleting the ten percent of the quatrains which have the biggest overlap with other quatrains until there is no overlap remaining.
+We refer to the resulting dataset (again) as QuaTrain.
+QuaTrain is split into train and dev sets using a ratio of 9:1; we do not keep a test set since no held-out human data is needed for generation or evaluation.
+Further, as some models used in this work are designed to process sonnets and/or limerick data, we create pseudo sonnets for them, denoted as SonNet.
+Specifically, for each sonnet, we randomly draw three quatrains and one couplet from the corresponding data split of QuaTrain, ensuring that each comes from a different original quatrain.
+Table 1 provides the data sizes.
+| Class | Model | Smaller | Larger | Lang |
+| Poetry- specific | DeepSpeare | - | - | de/en |
+| SA | - | - | de/en | |
+| Unconditioned / Conditioned LLMs | ByGPT5 | 140m | 290m | de/en |
+| GPT2 | 117m | 774m | de/en | |
+| GPTNeo | 125m | 1.3b | en | |
+| LLaMA2 | 7b | 13b | de/en | |
+| LLaMA3 | 8b | de/en | ||
+Models
+We use 2 different model classes:
+Poetry-specific Models: We select two models that integrate LSTM language models with additional components to generate quatrains with rhymes.
+DeepSpeare (Lau et al., 2018) utilizes a pentameter model to learn iambic meter and a rhyme model to distinguish between rhyming and non-rhyming words.
+Structured Adversary (SA) (Jhamtani et al., 2019) learns to rhyme in an adversarial setup, where a language model aims to generate poems misclassified by the discriminator, while a discriminator is trained to differentiate between generated and real poems.
+Both models can take sonnets as input during training and output quatrains during inference.
+For more detailed model descriptions, see Appendix A.1.
+General Purpose LLMs: We consider several decoder-only transformer-based models, encompassing both (sub)word- and character-level models, as well as older and very recent models.
+We choose two model families from the GPT series, GPT2 (Radford et al., 2019) and GPTNeo (Black et al., 2022) (a replicated version of GPT3 by EleutherAI222https://www.eleuther.ai/), two from the LLaMA series, LLaMA2 (Touvron et al., 2023) and LLaMA3 (AI@Meta, 2024), and the character-level ByGPT5 (Belouadi and Eger, 2023).
+Except for LLaMA3, we consider one smaller and one larger variant within each model family based on model size.
+We train each model in both unconditioned and conditioned manners, with rhymes and meters exposed during training in the latter case.
+We encode styles with special tokens during training and allow the models to predict the styles autonomously during inference.
+For all LLMs, we employ consistent decoding strategies for generation: we use the default settings of the LLaMA2 chat models on Hugging Face333https://huggingface.co/spaces/huggingface-projects/llama-2-7b-chat but limit the number of newly generated tokens to 100 for the word-level models and 300 for the character-level ByGPT5 models.
+We end up with a total of 36 models for German and English, categorized into three groups: 1) poetry specific LSTM-based models, 2) unconditioned LLMs, and 3) conditioned LLMs, as summarized in Table 2.
+SonNet is used for training 1), while QuaTrain is used for 2) and 3), separately for each language.
+We train all models using early stopping based on the perplexity/loss observed in the dev sets (details see Appendix A.2), as overfitting may negatively bias certain metrics like memorization rates.
+To distinguish between the different sizes and training manners of the LLMs, we use the following notation: a subscript of S/L indicates whether it is a smaller/larger version, and a superscript of “con” stands for conditioned training.
 E.g., GPT2 and GPT2 represent the unconditioned and conditioned trained GPT2 small models, respectively.
-Table 4 (appendix) provides an overview of all models belonging to the second and third category (transformer-based LLMs).
-5.3 Sampling
-From each model class, we randomly draw 1000 generated poems.
-Whenever wo do a direct comparison between training and generated data (e.g.
+| Memorization () | Length () | Rhyme () | ||||||
+| DE | EN | DE | EN | DE | EN | |||
+| Couplet | Verse | Couplet | Verse | |||||
+| Poetry-specific | 0.0000 | 0.006 | 0.0000 | 0.0046 | 0.752 | 0.745 | 0.992 | 0.825 |
+| Character-level | 0.0000 | 0.010 | 0.0000 | 0.0087 | 0.815 | 0.813 | 0.893 | 0.895 |
+| Word-level | 0.0476 | 0.048 | 0.0005 | 0.0309 | 0.686 | 0.700 | 1.057 | 0.852 |
+| Unconditioned | 0.0003 | 0.045 | 0.0006 | 0.0324 | 0.686 | 0.681 | 1.107 | 0.937 |
+| Conditioned | 0.0004 | 0.028 | 0.0002 | 0.0194 | 0.760 | 0.769 | 0.913 | 0.785 |
+| Larger | 0.0005 | 0.037 | 0.0005 | 0.0290 | 0.713 | 0.705 | 1.111 | 0.861 |
+| Smaller | 0.0003 | 0.039 | 0.0003 | 0.0237 | 0.726 | 0.756 | 0.931 | 0.890 |
+| Lexical () | Semantic () | |||||||||
+| DE | EN | DE | EN | |||||||
+| ATTR | MATTR | MTLD | ATTR | MATTR | MTLD | Within | Across | Within | Across | |
+| Poetry-specific | 0.928 | 0.895 | 162.8 | 0.890 | 0.863 | 126.0 | 0.577 | 0.669 | 0.509 | 0.601 |
+| Character-level | 0.915 | 0.886 | 166.7 | 0.837 | 0.818 | 83.4 | 0.582 | 0.678 | 0.522 | 0.610 |
+| Word-level | 0.922 | 0.874 | 114.7 | 0.871 | 0.835 | 82.7 | 0.629 | 0.693 | 0.587 | 0.634 |
+| Unconditioned | 0.919 | 0.875 | 125.9 | 0.854 | 0.818 | 75.2 | 0.613 | 0.688 | 0.580 | 0.632 |
+| Conditioned | 0.921 | 0.880 | 133.2 | 0.873 | 0.845 | 90.6 | 0.619 | 0.688 | 0.571 | 0.627 |
+| Larger | 0.932 | 0.890 | 143.9 | 0.873 | 0.837 | 84.1 | 0.613 | 0.689 | 0.571 | 0.626 |
+| Smaller | 0.902 | 0.861 | 115.6 | 0.839 | 0.814 | 74.3 | 0.623 | 0.688 | 0.577 | 0.631 |
+5 Evaluation
+We first report the results of diversity evaluation in §5.1, which is our main focus, followed by an examination of the relationship between diversity and overall quality through human evaluation in §5.2.
+5.1 Diversity Evaluation
+From each model, we randomly draw 1000 generated poems.
+Whenever we do a direct comparison between training and generated data (e.g.
 when comparing lexical diversity), we randomly draw 10 samples of size 1000 (matching the sample size) from the train set and use mean results as representatives.
 We deploy this strategy to mitigate the large discrepancy in size between human data and generated poems.
-For all LLMs, we employ consistent decoding strategies for generation: we use the default settings of the LLaMA2 chat models on Hugging Face111https://huggingface.co/spaces/huggingface-projects/llama-2-7b-chat but limit the number of newly generated tokens to 100 for the token-level models or 300 for the character-level ByGPT5 models.44todo: 4YC: @Hannes, do you have different sampling strategies for bygpt5?
-H: Other than the maximum length it’s the same.
-In Section 7, we briefly discuss the effects of various decoding strategies.
-6 Experiments and Results
 We first investigate structural properties of the generated poems (repetition of instances on a surface level, length distributions, rhyming), then consider lexical and semantic properties.
+After discussing each dimension of diversity, we provide a brief summary that generalizes across different model classes (e.g., poetry-specific vs.
+style conditioned vs.
+unconditioned, character- vs.
+word-level, larger vs.
+smaller).
+These summaries are based on Table 3.
 Memorization
-On quatrain-level, no model exhibits memorization, with all obtaining a memorization rate of 0.
+| DE | EN | |||
+| verse | couplet | verse | couplet | |
+| DeepSpeare | 0.83% | 0.83% | ||
+| SA | 0.40% | 0.10% | ||
+| ByGPT5 | 1.30%∗ | 1.23%∗ | ||
+| ByGPT5 | 1.23% | 0.93% | ||
+| GPT2 | 6.85% | 0.10% | 3.90% | 0.10% |
+| GPT2 | 8.70%∗ | 0.10% | 4.03%∗ | 0.10% |
+| GPTNeo | - | 5.60%∗ | 0.05% | |
+| GPTNeo | - | 4.73% | 0.10%∗ | |
+| LLaMA2 | 4.65% | 3.45%∗ | 0.05%∗ | |
+| LLaMA2 | 5.45%∗ | 2.48% | ||
+| LLaMA3 | 3.60% | 2.88% | 0.05% | |
+| ByGPT5 | 0.90%∗ | 0.58% | ||
+| ByGPT5 | 0.68% | 0.75%∗ | ||
+| GPT2 | 4.38% | 0.15%∗ | 2.33%∗ | 0.10%∗ |
+| GPT2 | 6.90%∗ | 0.10% | 2.03% | |
+| GPTNeo | - | 3.88%∗ | 0.05%∗ | |
+| GPTNeo | - | 3.50% | ||
+| LLaMA2 | 4.03%∗ | 0.05%∗ | 2.23%∗ | |
+| LLaMA2 | 0.70% | 0.55% | ||
+| LLaMA3 | 2.33% | 1.65% | ||
+Table 4 showcases the couplet- and verse level memorization rates. Since all models exhibit zero memorization rates on quatrain-level, we omit them in the table.
 Considering couplet-level memorization, 23 out of 36 models show zero memorization, while 13 models display scores between 0.05% and 0.15%.
@@
 When comparing models of the same architecture and training methods but varying sizes, differences are found in 6 out of 14 cases.
-In 5 cases, larger models have 0.05%-0.10% higher absolute memorization scores than their smaller counterparts (the English GPT2con and LLaMA2con models, and the German GPT2con, GPTNeocon, LLaMA2 models); the only exception is the English GPTNeo models, where the smaller one has a 0.05% higher memorization rate.
-On the other hand, there is no distinct boundary between unconditioned and conditioned trained models: in the 6 cases where discrepancies in memorization rates exist, conditioned models outperform their unconditioned ones three times, and vice versa, with absolute differences of 0.05%-0.1%.
+In 5 cases, larger models have 0.05%-0.10% higher absolute memorization scores than their smaller counterparts (the German GPT2con and LLaMA2con models, and the English GPT2con, GPTNeocon, LLaMA2 models); the only exception is the English GPTNeo models, where the smaller one has a 0.05% higher memorization rate.
+On the other hand, conditioned models mostly outperform their unconditioned counterparts: in 4 out of 6 cases where discrepancies in memorization rates exist, the conditioned ones exhibit lower memorization rates, with absolute declines of 0.05%-0.10%.
 In the verse-level evaluation, the poetry-specific models perform best overall (0.4%-0.83% for German and 0.1%-0.83% for English), followed by the ByGPT5 models (0.68%-1.3% for German and 0.58%-1.23% for English).
+SA is the best individual model, obtaining memorization rates of 0.4% for German and 0.1% for English.
 Again, GPT2 is worst for German, exhibiting memorization rates of 4.38%-8.7%, whereas, for English, GPTNeo exhibits the highest rates, ranging from 3.5%-5.6%.
 Concerning different model sizes, we again see that larger models memorize more than their smaller counterparts: in 9 out of 14 cases, larger models show higher memorization rates, with an average absolute increase of 0.15%.
-Furthermore, unlike in the couplet-level evaluation, here, each unconditioned model exhibits a strictly higher memorization rate compared to its conditioned counterpart, with an absolute increase of 1.47% on average.
-In summary, with respect to memorization: (1) Among model families, the poetry-specific and character-level models are the most diverse.
+Here, each conditioned model exhibits a strictly lower memorization rate compared to its unconditioned counterpart, with an absolute decrease of 1.47% on average.
+Overall: (1) No models exhibit severe memorization issues, such as copying entire poems or large portions of poem snippets from the training data.
+In terms of memorization, (2) among model groups, the poetry-specific and character-level models are more diverse; SA is the best individual one.
 (3) Larger models are less diverse compared to their smaller versions.
@@
 The poems generated by German LLaMA2 and English GPTNeo are too short and not diverse enough (in terms of standard deviation).
-The poetry-specific DeepSpeare models do not match the human distribution very well either, with intersection values of 0.63 and 0.57 for German and English respectively.
-Here too, poems lengths are too small and not diverse enough.
+The poetry-specific DeepSpeare models do not match the human distribution very well either, with intersection values of 0.63 and 0.57 for German and English, respectively.
+Here, too, poem lengths are too short and not diverse enough.
 Conditioned models seem to fit the training data better across the board, the only exceptions being German ByGPT5 and English LLaMA2.
-Figure 1 illustrates the length distribution of human written poems, SA and GPTNeo for English.
+Figure 4 (appendix) illustrates the length distribution of human written poems, SA and GPTNeo for English.
+Overall, regarding the alignment with human distributions: (1) Character-level ByGPT5 models generally align best with human data, followed by poetry-specific models; nevertheless, the poetry-specific SA is the top individual model.
+(2) Style-conditional models outperform the unconditioned trained ones.
+(3) Smaller models demonstrate a better fit than the larger ones.
 Rhyme
-Figure 2 (a) shows the distributions of rhyme schemes in our human training datasets for German; Figure 3 (a) shows the corresponding distribution for English.
-55todo: 5SE: Most rhymes in the training data are classified as real rhymes.
+Figures 2 (a) and 3 (a) show the distributions of rhyme schemes in our human training datasets for German and English, respectively.
 For both languages, less than 15% of all quatrains in training do not rhyme at all (rhyme scheme ABCD).
@@
 German has a higher proportion of ABAB (above 35%), while English has ABAB and AABB in roughly equal proportions (25%).
-In Table 6 (appendix), we report the entropy of all rhyme distributions as well the distance between the human distribution and model distributions, measured in cross-entropy and KL divergence.
-The best, worst and an average model, in terms of cross-entropy, are shown in Figures 2 and 3.
+Table 8 (appendix) reports the entropy of all rhyme distributions and the distance between the human distribution and model distributions, measured in KL divergence.
+The best, worst and an average model, in terms of KL divergence, are shown in Figures 2 and 3.
 Poetry-specific models: Figure 5 (appendix) shows the distributional plots for DeepSpeare and SA.
@@
 The three dominating patterns are AABB, ABAB, and ABBA which (only) partially agrees with the dominating patterns in the human data.
-Nonetheless, DeepSpeare has the best fit of all models in terms of KL divergence and almost the best fit in terms of cross-entropy.
+Nonetheless, DeepSpeare has the best fit of all models in terms of KL divergence, ranking first for German and second for English.
 SA has a much worse fit and produces considerably too many ABCD patterns (close to or above 30% in both languages).
@@
 Style-conditioned LLMs are shown in Figures 8 and 9 (appendix).
-In general, this model class matches the human distribution closest in terms of KL divergence and cross-entropy.
-LLAMA3 has a particularly good fit in English, but it is noteworthy that it produces too many ABAB rhyme schemes, which may indicate a skewness due to its pre-training data.
-No model produces a lot of AABB rhyme pattern which abound in our human training data.
+In general, this model class matches the human distribution closest in terms of KL divergence.
+However, no model produces a lot of AABB rhyme pattern which abound in our human training data.
 Across all models in this class, the fit to the human data is still mediocre at best.
 Overall, most models have clearly higher ABCD rhyming schemes than the human data, thus are underdiverse concerning rhyming.
-The best model class are style-conditioned LLMs (especially LLAMA3), however the poetry-specific DeepSpeare model can be considered the best individual model in terms of matching the human rhyme distribution.
-The character-level ByGPT5 models perform worse than word-level models without style-conditioning, but with style-conditioning they outperform the word-level models in terms of match with human rhyme distribution.
+(1) Conditioned models very clearly outperform unconditioned models and (2) character-level and poetry-specifc models are clearly better than word-level models in terms of matching the human rhyme distribution.
+(3) There is no clear size effect.
 | Model | ATTR (%) | MATTR (%) | MTLD |
@@
 | GPT2 | 89.2 / 84.0 | 84.2 / 81.9 | 82.0 / 70.3 |
-GPT2 | 94.2 / 88.0 | 90.0 / 90.7 | 137.4 / 90.7 |
+| GPT2 | 94.2 / 88.0 | 90.0 / 85.3 | 137.4 / 90.7 |
 | GPTNeo | - / 83.1 | - / 80.2 | - / 61.2 |
@@
 Table 5 shows the lexical diversity results for English and German.
-For local diversity (ATTR), most of the models are close to the diversity in human-written poems, with the traditional models (DS, SA) and the Llama exceeding the ATTR values of human-written poems.
-For English, the least locally diverse poems are generated by GPT2-small, in the un/conditioned case respectively.
-For German, the least locally diverse models is GPT2neo-small, in the un/conditioned case respectively.
+For local diversity (ATTR), most of the models are close to the diversity in human-written poems, with the traditional models (DeepSpeare, SA) and the LLaMA exceeding the ATTR values of human-written poems.
+For German, the least locally diverse poems are generated by GPT2, in the un/conditioned case, respectively.
+For English, the least locally diverse models is GPTNeo, in the un/conditioned case, respectively.
 The global diversity metrics (MATTR, MTLD) show different trends than ATTR, though.
-The MATTR metric suggests that most models do not generally achieve the level of diversity found in human poems: in English, only SA matches and slightly exceeds human diversity, in German, only the conditioned Llama3-large model exceeds human diversity.
+The MATTR metric suggests that most models do not generally achieve the level of diversity found in human poems: in English, only SA matches and slightly exceeds human diversity, in German, only the LLaMA2 and LLaMA3 model exceeds human diversity.
 According to the MTLD metric, all models generate severely under-diverse output at the sample level.
-Here, the best model in English and German is , but even does not come close to the human level of global diversity.
+Here, the best model in English and German is SA, but even SA does not come close to the human level of global diversity.
 According to MTLD, style-conditioned LLMs consistently outperform their non-conditioned counterparts, with the English LlaMA2 models being the only exceptions here.
-††todo: include barplots in the appendix?
 Moreover, we observe that model size affects all three lexical diversity metrics, whereby larger models are more diverse than their smaller counterparts.
-††todo: produce a barplot that shows the effect of size?
-The effect of size is most pronounced for GPT2 in both languages, where ATTR, MATTR and MTLD substantially improve from the small to the larger model variant.
-It also noteworthy, though, that the more classical models, and , generally perform on par with recent transformer models and sometimes even outperform them, as in the case of for global diversity.
-This shows that unconditional LLMs avoid repetitions at a local level whereas, at the sample level, they generate poems that are lexically much more similar to each other than poems within the human sample.
-††todo: explain the differences between MATTR and MTLD
+The effect of size is most pronounced for GPT2, where ATTR, MATTR and MTLD substantially improve from the small to the larger model variant.
+Generally, the MTLD results suggest more pronounced differences between models as well as humans and models than MATTR.
+Overall, in terms of lexical diversity, (1) neural models match human performance at the local level but fall short at the global level.
+(2) Poetry-specific models outperform other model classes, while character-level LLMs are most deficient (except for MTLD).
+(3) Conditional training is beneficial.
+(4) Larger models perform better.
 Semantic Similarity
@@
 | LLaMA2 | 62.3 / 58.0 | 68.9 / 62.9 |
-LLaMA2 | 61.2 / 58.4 | 69.1 / 63.8 |
+| LLaMA3 | 61.2 / 58.4 | 69.1 / 63.8 |
 | ByGPT5 | 58.4 / 52.2 | 67.7 / 60.8 |
@@
 | LLaMA2 | 63.3 / 58.5 | 69.5 / 62.9 |
-LLaMA3 | 59.6 / 58.2 | 68 / 62.3 |
+| LLaMA3 | 59.6 / 58.2 | 68.0 / 62.3 |
 Table 6 presents results for the semantic (cosine) similarity of quatrains: (i) within human and model-generated samples, and (ii) across generated samples and the human data.
-These results generally confirm the trends we found for global lexical diversity discussed above.
 None of the models generates a sample of poems with a within-sample diversity as low as the human with-sample diversity.
-SA is the model that achieves the lowest within-sample similarity and the lowest across-sample similarity, suggesting that it deviates most from the patterns in the human training data.
-Note that SA also achieved the best results in global lexical diversity (MATTR and MTLD in Table 2).††todo: SZ: I am not entirely sure about the across-sample similarities: is it necessarily a good thing if the models have low similarity with the human sample?.
-Moreover, the results on semantic similarity confirm the trends we observed with model size for lexical diversity, but disconfirm the trends for the effect of conditioning.
-Thus, we do not see a consistent trend for conditioned models generating samples with lower similarity/higher diversity.
-For model size, on the other hand, we observe a general trend towards larger models outperforming their smaller counterparts.
-7 Discussion
-Sampling/Searching
-We deploy various decoding strategies to determine to what extent these can alter the various aspects of diversity in the generated poems.
-We use different combinations of temperature-based sampling (ackley1985learning), Nucleus sampling (Top-p) (holtzman2019curious) and Top-k sampling (fan-etal-2018-hierarchical) as sampling strategies and further deploy two variants of contrastive search (su2022contrastive).
-Results indicate that the various techniques can only slightly increase diversity in one or more aspects.
-Moreover, aggressive sampling often leads to output degeneration, causing the models to (partially) repeat verses in a quatrain.
-We provide some examples in the appendix, see Tables 7 to 12 as well as Figures 10 to 15.
+SA is the model that achieves the lowest within-sample similarity and the lowest across-sample similarity.
+Overall, (1) poetry-specific models are most diverse regarding semantic similarity and word-level models are least diverse; (2) style-conditioning makes models slightly more diverse semantically; (3) larger models are also slightly more diverse.
 Which is the most diverse model?
-We have seen that unconditioned LLMs exhibit poor results with regard to different dimensions of diversity: they do not rhyme, are lexically underdiverse and do not show sufficient semantic variation.
+We have seen that unconditioned LLMs exhibit poor results across various dimensions of diversity: they often do not rhyme, are lexically underdiverse and do not show sufficient semantic variation.
 However, character-level models are more diverse than word level models.
-Style-conditioned models perform better regarding rhyming, semantic variation, and lexical variation but word level style-conditioned models are prone to severe memorization from the training data, in particular when the model is large and the training set is small.
+Style-conditioned models perform better regarding memorization, rhyming, and lexical variation, while deviating less from human poems according to the distribution match of length and rhymes.
+On the other hand, larger LLMs often outperform their smaller counterparts in semantic and lexical diversity, but they also tend to memorize more from the training data.
 Character-level style-conditioned LLMs produce overall best diversity results and do not deteriorate as a function of model/training data size.
-In terms of diversity, poetry-specific Deepspeare performs similar as character-level LLMs but requires more modeling effort from human experts (e.g., in developing rhyming components).
+In Appendix A.3, we calculate the average ranks of the models across all 5 dimensions, finding that indeed, for both languages, the conditioned trained ByGPT5 models perform overall best among all models, ranking as the first and second places for German and the first and third places for English.
+In terms of diversity, poetry-specific SA and DeepSpeare overall lag only slightly behind character-level LLMs but require more modeling effort from human experts (e.g., in developing rhyming components).
+The largest word-level LLMs explored in this work, LLaMA2 and LLaMA3, generally perform best among the word-level models; however, they do not exhibit superiority over the style-conditioned character-level models and poetry-specific models as well.
+We also compute Pearson’s correlations between ranks for different dimensions.
+For German, the highest correlation is between semantic diversity and memorization (0.842), followed by the two moderate to high correlations: 0.526 (semantic vs.
+lexical) and 0.518 (memorization vs.
+rhyme).
+Two pairs show moderate correlations: 0.480 (semantics vs.
+length) and 0.404 (memorization vs.
+rhyme).
+The remaining pairs exhibit weak positive or negative correlations, with absolute values between 0.051 and 0.228.
+For English, no pairs exhibit high correlations.
+Two pairs show moderate to high correlations: 0.628 and 0.635 (memorization vs.
+semantics/length).
+Three pairs demonstrate moderate correlations, ranging from 0.307 to 0.357 (semantics vs.
+lexical/length and memorization vs.
+length).
+The others show weak correlations, with absolute values between 0.024 and 0.267.
+Concretely, these sometimes low correlations are mirrored in the different ranks models have across different dimensions: for example, SA is almost as diverse as the human training data regarding semantics and length, but provides one of the worst fits regarding rhyming.
+This indicates that most current models face a tradeoff for different diversity dimensions.
+5.2 Quality Evaluation
+Diversity in model outputs could sometimes result from low coherence or a lack of meaningful content.
+To investigate whether this is the case, we conducted a small-scale human evaluation of the overall quality of quatrains, focusing specifically on coherence and semantics (punctuation was omitted here, as it was also excluded during the diversity evaluation).
+In this evaluation, we compared 60 outputs across 5 systems (12 outputs per system) for each language, including human-written quatrains, and the outputs of the winning models in overall, lexical, semantic, and rhyme diversity (as presented in Tables 9 and 10 in the appendix).
+We created 15 annotation instances; in each instance, an annotator was given 4 quatrains and asked to select both the best and the worst among them.
+As Table 11 in the appendix displays, for German, human quatrains are clearly preferred (they were chosen as the best 12 times and the worst 0 times).
+The best automatic system is the overall winning ByGPT5 model (best 2 times; worst 1 time); SA is the worst (worst 8 times).
+For English, the lexical winning LLaMA3 model is the best in terms of coherence (best 6 times; worst 0 times), followed by the rhyme winning GPTNEO model (best 5 times; worst 0 times); SA is again the worst (worst 11 times).
+However, we noted that our evaluator was a native speaker of German but not English and said that the German evaluation was much easier for him.
+The older SA model appears to have higher diversity at the cost of quatrain quality.
+However, overall, we conclude that more diverse models also seem to be qualitatively better — this does not have to be a causal/strong relationship, however, especially for the newer LLMs.
+Tables 12 and 13 in the appendix present 10 sample quatrains selected as the best in our human evaluation, including both system-generated and human-written ones.
 6 Conclusion
-To date, evaluation of automatic poetry generation has almost exclusively focused on human evaluation and forms of the Turing test.
-Our work shows that an automatic assessment of the diversity of generated poems covers an important blind spot of existing studies.
-Our evaluations shed light on the fact that none of the state-of-the-art poetry generators is able to match the level of diversity in human poems, confirming previous evaluations of diversity in other NLG tasks ippolito-etal-2019-comparison; schuz-etal-2021-diversity; stasaski-hearst-2022-semantic.
-Our study also adds a new dimensions to previous work on diversity, by showing that diversity on the level of rhyming is particularly hard to achieve for neural generators and interacts with other dimensions of diversity in poetry generation, i.e., style conditioned LLMs do not only achieve a better match with human rhyme and length distributions, but also higher lexical and semantic diversity .
-We also find that memorization — a general and widely discussed limitation of LLMs (carlini2021extracting) — is a potential issue in poetry generation, especially for certain combinations of model sizes and finetuning schemes, complementing existing studies in this area mireshghallah-etal-2022-empirical.
-We release all code upon acceptance.
+Our work is the first and most comprehensive automatic evaluation of poetry diversity, yielding several interesting observations: for example, we find that style-conditioning enhances virtually all measures of diversity and that character-level modeling also increases diversity, including reducing memorization.
+Our evaluations also shed light on the fact that none of the state-of-the-art poetry generators is able to match the level of diversity in human poems.
+Thus, we find overall that an automatic assessment of the diversity of generated poems covers an important blind spot of existing studies.
+Future work should aim for more diverse automatic poetry generation systems as a prerequisite of general computational creativity.
 Limitations
@@
 word), further aspects — and in particular pre-training — may be worth investigating in future work.
-Generally, our work is concerned with the evaluation of NLG systems; evaluation methods and evaluation metrics (zhao-etal-2019-moverscore; Zhang2020BERTScore; NEURIPS2021_e4d2b6e6; chen_menli:2023; peyrard-etal-2021-better) are a well-known and notorious issue in this research field.
-While a lot of recent work has aimed at improving common practices in human evaluation (belz-etal-2023-missing) or advancing the study of metrics for quality or fluency of NLG outputs, the evaluation of diversity is comparatively under-researched.
+Due to the hardware constraints and time limitations, we did not run experiments multiple times to take the averages or optimize the training hyperparameters, which may have introduced a degree of randomness in our results.
+For example, in our initial experiments, we trained GPT2 models with a slightly different setting.
+Compared to the GPT2 models we mainly reported, these models behave slightly differently.
+E.g., they exhibit better lexical diversity, as shown by an increase in ATTR from 0.87 to 0.89, MATTR from 0.84 to 0.86, and MTLD from 88 to 101 on average.
+Similarly, they are also more diverse according to the semantic similarity metrics, which are on average 0.02-0.03 lower.
+In contrast, these models perform worse in rhyming; they have a 10% lower chance of producing rhymed quatrains, and their rhyme distributions are more distant from human distributions (0.27 higher KL divergence).
+Despite these differences, our findings are generally robust as we report averages over model classes in our analysis.
+For the same reason, we did not select the largest versions of these models; nevertheless, our evaluation already shows prominent differences in diversity across model sizes.
+Further, we note that our trained LLMs occasionally do not generate texts in the form of a quatrain (i.e., 4 verses).
+These outputs were excluded from the analysis, though such cases are rare (1.5% on average).
+Ethics Statement
+All the datasets, models, and code used in this work will be made publicly available.
+We have not collected private or sensitive data and have only used language models with free access, such that our experiments can be fully replicated by anyone.
+Generally, our work is concerned with the evaluation of NLG systems; evaluation methods and evaluation metrics (Zhao et al., 2019; Zhang et al., 2020; Peyrard et al., 2021; Yuan et al., 2021; Chen et al., 2022; Chen and Eger, 2023; Leiter et al., 2023) are a well-known and notorious issue in this research field.
+While a lot of recent work has aimed at improving common practices in human evaluation (Belz et al., 2023) or advancing the study of metrics for quality or fluency of NLG outputs, the evaluation of diversity is comparatively under-researched.
 In this work, we aimed at providing a range of metrics assessing different aspects of diversity, but could not cover all potentially interesting ways of measuring diversity.
@@
 Future work could also consider more (diverse) languages and other genres and datasets for poetry.
-Ethics Statement
-Often, the discussion of creative AI systems in public discourse is surrounded by misconceptions, hypes and even myths (veale2012exploding).
-Our work contributes to a careful operationalization and objective assessment of the creative capbalities of AI systems in the area of poetry generation.
-All the datasets, models and code used in this work are publicly available or will be made available upon publication.
-We have not collected private or sensitive data and have only used language models with free access, such that our experiments can be fully replicated by anyone.
 Appendix A Appendix
-Model | EN | DE |
-GPT2-small | 124M | 124M |
-GPT2-large | 774M | 774M |
-GPTneo-small | 125M | - |
-GPTneo-xl | 1.3B | - |
-byGPT5-base | 140M | 140M |
-byGPT5-medium | 290M | 290M |
+A.1 DeepSpeare and SA
+Deepspeare Lau et al.
+(2018) is specifically designed for poetry generation.
+Its core architecture consists of an LSTM language model, a pentameter model (specifically designed to learn iambic meter) and a rhyme model.
+During training, it takes sonnets as input data (three quatrains followed by a couplet) but ultimately processes the contained quatrains by splitting any given sonnet.
+The rhyme model processes ending words of quatrain verses and uses a margin-based loss to discriminate between rhyming and non-rhyming words.
+It is not limited to specific rhyme patterns but assumes that rhymes exist in the data.
+At inference time, Deepspeare generates quatrains.
+Structured Adversary.
+Like Deepspeare, Structured Adversary (SA) Jhamtani et al.
+(2019) incorporates different components: an LSTM language model and a discriminator used to decide whether line endings are typical for poetry.
+Both components are organized in an adversarial setup, where the language model acts as a generator, trying to generate poems that are misclassified by the discriminator, while the discriminator is trained to distinguish generated poems from real ones.
+SA is trained with sonnets as input data.
+At inference time, it generates quatrains.
+A.2 Training
+DeepSpeare
+DeepSpeare (Lau et al., 2018) leverages pretrained static word vectors.
+We use QuaTrain and SonNet to train our own Word2vec embeddings (Mikolov et al., 2013) and the final sonnet models respectively.
+For the sonnet model training, we use a batch size of 128 and apply early stopping with a patience of 5 epochs; default settings are maintained for the other hyperparameters.
+SA
+We use the same word vectors and training data splits as for DeepSpeare.
+Training SA involves 1) pretraining the discriminator’s encoder using a publicly available pronouncing dictionary; 2) training the LM component; 3) training a final aggregated model in a generative adversarial setup.
+We train the discriminators with a batch size of 128, the LMs with a batch size of 64, and the final sonnet models with a batch size of 128; here, we also implement early stopping with a patience of 5 epochs.
+Style-un/conditioned LLMs
+We train all LLMs for 50 epochs on our train set using the paged AdamW optimizer with a weight decay of 0.001, a learning rate of 4e-05, a cosine learning rate decay with a 3% warmup ratio, and early stopping with patience of 5 epochs.
+As we run experiments on GPUs with varying memory capacities ranging from 12GB to 80GB, and with models that drastically differ in size, to achieve as much consistency as possible, we either train models with a batch size of 128 or accumulate the batches to reach a size of 128.
+For LLaMA, we use 4-bit quantization and LORA (Hu et al., 2021); the corresponding parameters are list below:
+target modules: q_proj, v_proj, k_proj, o_proj, embedded_tokens
+lora alpha: 16
+lora dropout: 0.05
+r: 16
+A.3 Evaluation Results
 | model | |||||||
@@
 | en | LLaMA3 | 0.78 | 16 | 65 | 27.12 | 26 | 5.35 |
-A.1 Rhyme
-DE | EN | |||||
-Model | Entropy | KL Divergence | Cross Entropy | Entropy | KL Divergence | Cross Entropy |
-HUMAN | 2.90 | 0.00 | 2.90 | 3.10 | 0.00 | 3.10 |
-DeepSpeare | 2.97 | 0.55 | 3.52 | 3.16 | 0.48 | 3.64 |
-SA | 3.14 | 1.43 | 4.57 | 3.22 | 1.17 | 4.39 |
-ByGPT5 | 2.89 | 1.23 | 4.12 | 2.92 | 1.08 | 4.00 |
-ByGPT5 | 3.13 | 1.09 | 4.21 | 2.91 | 1.13 | 4.04 |
-GPT2 | 2.86 | 1.26 | 4.13 | 2.97 | 1.06 | 4.03 |
-GPT2 | 3.16 | 1.13 | 4.29 | 2.99 | 1.03 | 4.03 |
-GPTNeo | - | - | - | 2.80 | 1.18 | 3.99 |
-GPTNeo | - | - | - | 3.16 | 0.96 | 4.11 |
-LLaMA2 | 2.93 | 1.18 | 4.11 | 3.24 | 0.71 | 3.95 |
-LLaMA2 | 3.18 | 1.04 | 4.21 | 3.24 | 0.71 | 3.95 |
-LLaMA3 | 3.27 | 0.83 | 4.10 | 3.45 | 0.56 | 4.02 |
-ByGPT5 | 3.17 | 0.67 | 3.84 | 3.22 | 0.83 | 4.05 |
-ByGPT5 | 3.16 | 0.58 | 3.75 | 3.38 | 0.54 | 3.92 |
-GPT2 | 2.98 | 0.99 | 3.97 | 3.41 | 0.61 | 4.02 |
-GPT2 | 3.11 | 1.04 | 4.14 | 3.22 | 0.85 | 4.07 |
-GPTNeo | - | - | - | 3.43 | 0.45 | 3.89 |
-GPTNeo | - | - | - | 3.29 | 0.83 | 4.11 |
-LLaMA2 | 2.69 | 1.33 | 4.02 | 2.89 | 0.95 | 3.85 |
-LLaMA2 | 3.11 | 0.71 | 3.83 | 2.67 | 1.07 | 3.74 |
-LLaMA3 | 2.98 | 1.06 | 4.04 | 2.58 | 0.94 | 3.52 |
-Appendix B Decoding methods
-Method | GPT2-large | p-GPTneo-xl | p-byGPT5-medium |
-0.002 | 0.89 | 0.006 | |
-0.000 | 0.96 | 0.006 | |
-0.000 | 0.98 | 0.004 | |
-0.002 | 0.99 | 0.010 | |
-0.004 | 0.98 | 0.004 | |
-0.012 | 0.99 | 0.020 | |
-0.002 | 0.99 | 0.010 | |
-0.000 | 0.99 | 0.008 | |
-0.002 | 0.98 | 0.004 | |
-0.004 | 0.99 | 0.012 | |
-0.004 | 0.94 | 0.004 |
-Method | mean | std | min | max | |
-0.002 | 0.55 | 0.06 | 0.33 | 0.77 | |
-0.000 | 0.57 | 0.06 | 0.37 | 0.82 | |
-0.000 | 0.58 | 0.06 | 0.39 | 0.72 | |
-0.002 | 0.57 | 0.07 | 0.38 | 0.81 | |
-0.004 | 0.57 | 0.06 | 0.37 | 0.97 | |
-0.012 | 0.59 | 0.07 | 0.44 | 0.99 | |
-0.002 | 0.56 | 0.06 | 0.37 | 0.91 | |
-0.000 | 0.58 | 0.07 | 0.33 | 0.84 | |
-0.002 | 0.57 | 0.07 | 0.42 | 0.95 | |
-0.004 | 0.57 | 0.07 | 0.39 | 0.98 | |
-0.004 | 0.56 | 0.06 | 0.34 | 0.94 |
-Method | mean | std | min | max | h | |
-29.83 | 13.64 | 4 | 150 | 0.75 | 5.46 | |
-37.77 | 20.58 | 4 | 150 | 0.87 | 4.48 | |
-35.67 | 8.85 | 13 | 64 | 0.86 | 1.89 |
-Model | lang | real rhymes | repetitions |
-en | 0.72 | 0.00 | |
-en | 0.84 | 0.00 | |
-en | 0.44 | 0.08 | |
-en | 0.53 | 0.21 | |
-de | 0.57 | 0.00 | |
-de | 0.66 | 0.00 |
-method | real rhymes | repetitions |
-0.156 | 0.022 | |
-0.196 | 0.052 | |
-0.164 | 0.071 | |
-0.176 | 0.220 | |
-0.170 | 0.103 | |
-0.181 | 0.124 | |
-0.146 | 0.030 | |
-0.168 | 0.131 | |
-0.191 | 0.112 | |
-0.175 | 0.153 | |
-0.194 | 0.087 |
-Method | ATTR | MATTR | MTLD |
-0.804 | 0.794 | 64.679 | |
-0.757 | 0.744 | 44.942 | |
-0.716 | 0.717 | 36.561 | |
-0.537 | 0.537 | 17.297 | |
-0.720 | 0.690 | 33.151 | |
-0.419 | 0.461 | 14.730 | |
-0.586 | 0.622 | 21.757 | |
-0.611 | 0.588 | 21.005 | |
-0.766 | 0.754 | 46.107 | |
-0.642 | 0.637 | 24.289 | |
-0.797 | 0.779 | 58.490 |
+Length
+Rhyme
+Table 8 shows the entropy of the rhyme distributions in each sample as well as the distances of the distributions to that in the human data, measured by KL divergence.
+Figure 3 demonstrates the human rhyme distribution as well as the best, worst, and an average fit distributions in terms of KL divergence.
+Figures 5, 6/7, and 8/9 demonstrate the rhyme distributions for the poetry specific models, unconditioned and conditioned LLMs, respectively.
+| DE | EN | |||
+| Model | Entropy | KL Divergence | Entropy | KL Divergence |
+| HUMAN | 2.90 | 0.00 | 3.10 | 0.00 |
+| DeepSpeare | 2.97 | 0.55 | 3.16 | 0.48 |
+| SA | 3.14 | 1.43 | 3.22 | 1.17 |
+| ByGPT5 | 2.89 | 1.23 | 2.92 | 1.08 |
+| ByGPT5 | 3.13 | 1.09 | 2.91 | 1.13 |
+| GPT2 | 2.86 | 1.26 | 2.97 | 1.06 |
+| GPT2 | 3.16 | 1.13 | 2.99 | 1.03 |
+| GPTNeo | - | - | 2.80 | 1.18 |
+| GPTNeo | - | - | 3.16 | 0.96 |
+| LLaMA2 | 2.93 | 1.18 | 3.24 | 0.71 |
+| LLaMA2 | 3.18 | 1.04 | 3.24 | 0.71 |
+| LLaMA3 | 3.27 | 0.83 | 3.45 | 0.56 |
+| ByGPT5 | 3.17 | 0.67 | 3.22 | 0.83 |
+| ByGPT5 | 3.16 | 0.58 | 3.38 | 0.54 |
+| GPT2 | 2.98 | 0.99 | 3.41 | 0.61 |
+| GPT2 | 3.11 | 1.04 | 3.22 | 0.85 |
+| GPTNeo | - | - | 3.43 | 0.45 |
+| GPTNeo | - | - | 3.29 | 0.83 |
+| LLaMA2 | 2.69 | 1.33 | 2.89 | 0.95 |
+| LLaMA2 | 3.11 | 0.71 | 2.67 | 1.07 |
+| LLaMA3 | 2.98 | 1.06 | 2.58 | 0.94 |
+Best model
+We rank the models for each dimension and then average the ranks across the five dimensions to determine the overall rankings.
+For dimensions with multiple metrics, such as the three memorization metrics (due to different evaluation levels) and the three lexical metrics (measuring local or global lexical diversity), we first rank the models according to each metric and then average these ranks to represent that dimension.
+For dimensions primarily based on distributions, we use metrics that measure the distance/similarity of their distributions from human data: KL divergence for rhyme and histogram intersection for length.
+The results are shown in Table 9 and 10 for German and English respectively.
+| Language | Model | Size | Conditioned | semantic | lexical | length | rhyme | memorization | avg_rank |
+| de | BYGPT5 | L | TRUE | 2.0 | 4.0 | 5.0 | 3.0 | 1.7 | 3.1 |
+| de | BYGPT5 | S | TRUE | 3.5 | 6.0 | 4.0 | 2.0 | 1.3 | 3.4 |
+| de | SA | - | - | 1.0 | 2.7 | 1.0 | 16.0 | 2.0 | 4.5 |
+| de | DS | - | - | 5.0 | 10.3 | 12.0 | 1.0 | 1.0 | 5.9 |
+| de | BYGPT5 | S | FALSE | 6.0 | 11.0 | 2.0 | 10.0 | 2.7 | 6.3 |
+| de | BYGPT5 | L | FALSE | 4.0 | 8.3 | 6.0 | 13.0 | 3.0 | 6.9 |
+| de | LLAMA3 | - | FALSE | 9.5 | 6.3 | 9.0 | 5.0 | 6.0 | 7.2 |
+| de | LLAMA3 | - | TRUE | 6.5 | 7.3 | 8.0 | 9.0 | 5.7 | 7.3 |
+| de | LLAMA2 | S | TRUE | 13.5 | 13.0 | 3.0 | 4.0 | 4.0 | 7.5 |
+| de | GPT2 | L | TRUE | 12.5 | 4.7 | 7.0 | 6.0 | 8.3 | 7.7 |
+| de | LLAMA2 | L | FALSE | 9.5 | 2.7 | 16.0 | 12.0 | 5.3 | 9.1 |
+| de | LLAMA2 | S | FALSE | 8.0 | 10.0 | 15.0 | 8.0 | 5.0 | 9.2 |
+| de | GPT2 | L | FALSE | 14.0 | 5.7 | 10.0 | 14.0 | 8.7 | 10.5 |
+| de | GPT2 | S | TRUE | 15.0 | 15.0 | 11.0 | 7.0 | 6.3 | 10.9 |
+| de | LLAMA2 | L | TRUE | 12.5 | 13.0 | 13.0 | 15.0 | 8.0 | 12.3 |
+| de | GPT2 | S | FALSE | 13.5 | 16.0 | 14.0 | 11.0 | 7.7 | 12.4 |
+| Language | Model | Size | Conditioned | semantic | lexical | length | rhyme | memorization | avg_rank |
+| en | BYGPT5 | S | TRUE | 3.5 | 11.7 | 4.0 | 3.0 | 2.0 | 4.8 |
+| en | SA | - | - | 1.0 | 4.0 | 1.0 | 19.0 | 1.0 | 5.2 |
+| en | BYGPT5 | L | TRUE | 2.0 | 9.7 | 5.0 | 9.0 | 1.7 | 5.5 |
+| en | DS | - | - | 3.5 | 9.0 | 17.0 | 2.0 | 2.3 | 6.8 |
+| en | LLAMA2 | S | FALSE | 17.5 | 5.7 | 2.0 | 6.0 | 4.7 | 7.2 |
+| en | LLAMA3 | - | TRUE | 12.0 | 1.7 | 9.0 | 11.0 | 3.3 | 7.4 |
+| en | GPT2 | L | TRUE | 9.0 | 9.0 | 6.0 | 5.0 | 9.3 | 7.7 |
+| en | LLAMA2 | L | TRUE | 12.0 | 5.0 | 7.0 | 12.0 | 4.0 | 8.0 |
+| en | LLAMA2 | S | TRUE | 7.0 | 3.3 | 13.0 | 16.0 | 1.3 | 8.1 |
+| en | LLAMA3 | - | FALSE | 13.0 | 3.0 | 16.0 | 4.0 | 9.0 | 9.0 |
+| en | LLAMA2 | L | FALSE | 9.0 | 6.3 | 15.0 | 7.0 | 10.3 | 9.5 |
+| en | GPT2 | S | TRUE | 17.5 | 14.0 | 3.0 | 10.0 | 3.7 | 9.6 |
+| en | BYGPT5 | L | FALSE | 5.5 | 15.7 | 10.0 | 17.0 | 3.0 | 10.2 |
+| en | BYGPT5 | S | FALSE | 5.5 | 17.3 | 8.0 | 18.0 | 2.7 | 10.3 |
+| en | GPTNEO | L | TRUE | 13.5 | 13.0 | 19.0 | 1.0 | 10.0 | 11.3 |
+| en | GPTNEO | S | TRUE | 16.0 | 17.0 | 11.0 | 8.0 | 5.7 | 11.5 |
+| en | GPT2 | L | FALSE | 10.5 | 11.0 | 12.0 | 15.0 | 11.3 | 12.0 |
+| en | GPT2 | S | FALSE | 17.0 | 19.0 | 14.0 | 14.0 | 11.7 | 15.1 |
+| en | GPTNEO | S | FALSE | 17.5 | 20.0 | 18.0 | 13.0 | 12.0 | 16.1 |
+| en | GPTNEO | L | FALSE | 17.5 | 14.7 | 20.0 | 20.0 | 11.3 | 16.7 |
+| DE | EN | |||||||||
+| System | HUMAN | overall | semantic | lexical | rhyme | HUMAN | overall | semantic | lexical | rhyme |
+| (ByGPT5) | (SA) | (SA) | (DeepSpeare) | (ByGPT5) | (SA) | (LLaMA3) | (GPTNeo) | |||
+| Best | 12 | 2 | 0 | - | 1 | 3 | 1 | 0 | 6 | 5 |
+| Worst | 0 | 1 | 8 | - | 6 | 2 | 2 | 11 | 0 | 0 |
+| BWS | 0.8 | 0.07 | -0.53 | - | -0.33 | 0.07 | -0.07 | -0.73 | 0.4 | 0.33 |
+| Quatrain | System | ||||
+ByGPT5 | ||||
+ByGPT5 | ||||
+DeepSpeare | ||||
+HUMAN | ||||
+HUMAN |
+| Quatrain | System | ||||
+LLaMA3 | ||||
+GPTNeo | ||||
+LLaMA3 | ||||
+HUMAN | ||||
+HUMAN |